### v1

# Convexity on manifolds without focal points and spectrum of the Laplacian

Abstract.
Convexity on complete Riemannian manifolds has typically been studied under constraints on the sign of sectional curvatures. This is especially so in the study of the effect of convexity on the spectrum of the Laplacian. In this article, we use the language of Jacobi tensors to construct a strictly convex function with bounded gradient on complete simply connected manifolds without focal points having negative Ricci curvature. From this we obtain a lower bound for the growth of the volume of geodesic balls, if the manifold is also assumed to be Kähler. Our primary applications concern the spectrum of the Laplacian. We use the strictly convex function which we construct to show that the spectrum of the Laplacian is absolutely continuous on a class of asymptotically harmonic manifolds. In fact, we completely determine the spectrum of the Laplacian on simply connected noncompact nonflat harmonic manifolds without focal points. For asymptotically harmonic manifolds we also show that equality holds in Cheeger’s inequality for the bottom of the spectrum, where Cheeger’s constant is equal to the constant mean curvature of the horospheres. We show that the well-known description of the spectrum of the Laplacian on Riemannian symmetric spaces of noncompact type and arbitrary rank can also be recovered using these geometric methods.
1. Introduction
Convex functions have long played a fundamental role in Riemannian geometry. By the convexity of a continuous real-valued function on a Riemannian manifold we mean that it is convex in the usual sense along any geodesic. The mere existence or nonexistence of convex functions have geometric and topological implications. Nonconstant convex functions on complete Riemannian manifolds exist only on those with infinite volume [2, Proposition 2.2]. In particular, they don’t exist on compact Riemannian manifolds. There are profound structural results arising from the interaction of curvature and convexity. Under appropriate curvature conditions there are natural geometrically defined convex functions on a Riemannian manifold. On complete simply connected Riemannian manifolds of nonpositive sectional curvature, called Hadamard manifolds, the square of the distance function from a fixed but arbitrary point and the Busemann functions are examples of convex functions which play a crucial role in the geometry of these manifolds. While the former is a smooth convex function the latter is merely a $C^{2}$-convex function [12, §1.6.6]. In the nonnegative sectional curvature situation for instance, a key idea in the fundamental work [8] is the construction of suitable continuous convex functions related to Busemann functions which have appropriate additional properties.
A function which is at least $C^{2}$ is (strictly) convex if its hessian is positive (definite) semi-definite everywhere. If a Riemannian manifold admits a strictly convex function, then it has no compact immersed minimal submanifolds [20, §9.2 C]. In particular, it has no closed geodesics. Such a manifold also allows for a generalization of Cartan’s theorem [29] when this strictly convex function has a minimum: every compact subgroup of the isometry group has a common fixed point. The existence of strict convex functions also has consequences for the $\mathrm{L^{2}}$-spectrum of the Laplacian. The study of the spectrum of the Laplacian on complete Riemannian manifolds has a long, rich history and there are very many results, often with various additional constraints placed on the structure of the manifold and on sectional curvature. For a succint recent overview of some of these aspects, we refer to [9]. We content ourselves by mentioning a few results which use convexity in an essential manner to deduce information on the spectrum to put our work in context, with no pretensions to completeness.
The absolutely continuous part of the spectrum of the Laplacian on complete Riemannian manifolds supporting a sufficiently regular strictly convex function is nontrivial [28, Theorem 2] under certain boundedness conditions on the gradient and the Laplacian of the said function. On complete simply connected manifolds with nonpositive sectional curvature, [28, Theorem 1] exhibited a strictly convex function related to the square of the distance function satisfying the estimates required to ensure the nontriviality of the absolutely continuous spectrum. Moreover, it was also shown that in such a setting there are no eigenvalues in the spectrum. In the case of nonnegative sectional curvature, explicit convex functions have been used to show that there are no eigenfunctions under the additional assumptions on the codimension of the soul of the manifold being $1$ or having a quadratic decay of radial curvatures when the soul is just a point [14, Theorems 4.2, 4.3].
One of the key results of this article is the construction of a $C^{2}$-strictly convex function with bounded gradient on complete simply connected Riemannian manifolds without focal points having negative Ricci curvature (Theorem 3.1). A distinguishing feature here is that we do not make any assumptions on the sign of the sectional curvature. It is well known [18] that there are both compact and noncompact manifolds without focal points whose sectional curvatures change sign. Since the Ricci curvature is an average of sectional curvatures, the sign of the Ricci curvature is a weaker notion when compared to the sign of the sectional curvature. That this is true especially for negative Ricci curvature can be seen from the remarkable result [23, Theorem A] that any smooth manifold of dimension at least three admits a complete Riemannian metric with negative Ricci curvature. Interesting classes of manifolds satisfying our assumptions include Riemannian symmetric spaces of noncompact type as well as simply connected noncompact nonflat harmonic manifolds without focal points. Harmonic manifolds are Einstein [21] i.e, $Ric_{g}=\kappa g$, and have horospheres, which are level sets of Busemann functions, of constant mean curvature $h\geq 0$. If the Ricci curvature $\kappa>0$, then the harmonic manifold is necessarily a compact rank one symmetric space and if $\kappa=0$, then it is necessarily Euclidean. All other simply connected harmonic manifolds have negative Ricci curvature. We remark here that $h>0$ for such manifolds, since a harmonic manifold is Euclidean when $h=0$ [21, §5, Theorem 19].
For the construction of the strictly convex function we make use of Jacobi tensors introduced in [13] and its relation to the Hessian of Busemann functions for manifolds of continuous asymptote. In particular, for manifolds without focal points these Hessians can be seen to be nonnegative. So each Busemann function is convex but not strictly convex since its Hessian vanishes in the direction of the gradient. Motivated by the work done in the context of symmetric spaces of noncompact type in [10], we then average the Busemann functions over the boundary of the geodesic compactification for manifolds without focal points. Here we use [16] for the geometry at infinity of such manifolds. Under the assumption of negative Ricci curvature, a strictly convex $C^{2}$-function with bounded gradient is then obtained. As an immediate consequence, we have that the manifolds under consideration admit smooth strictly convex functions which also has bounded gradient. This follows from the approximation result of Green and Wu [17, Theorem 2(a)]. It is known [26] that the smooth approximation result fails for functions which are merely convex.
Our next application is in the context of manifolds which are, in addition, Kähler. A complex manifold with a complex structure $J$ and a compatible Riemannian metric $g$ is called Kähler if the associated alternating two-form defined by $\omega(X,Y):=g(JX,Y)$ is symplectic. Given any two of these compatible structures the third one can be recovered. ${\mathbb{C}}^{n}$ with the standard metric, bounded domains in ${\mathbb{C}}^{n}$ with the Bergmann metric, Hermitian symmetric spaces with the invariant metric, complex projective spaces with the Fubini-Study metric are some illustrative examples of Kähler manifolds. $C^{2}$-strictly convex functions on Kähler manifolds can be seen to be strictly plurisubharmonic functions. In our context, therefore, we get a smooth strictly plurisubharmonic function with bounded gradient. From this we derive a lower bound for the growth of the volume of metric balls on such manifolds following [24, Lemma 4.2 (i)]. The growth of the volume of the metric balls has been usually studied with lower bounds on Ricci curvature and upper bounds on sectional curvatures [7, §III.4].
As principal applications of Theorem 3.1, we describe the spectrum of the Laplacian for certain classes of manifolds having horospheres of constant mean curvature $h$. Complete simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$ are known as asymptotically harmonic manifolds [21]. On manifolds as in Theorem 3.1, which are in addition asymptotically harmonic, we observe that the strictly convex function that we construct also has bounded Laplacian and proceed to show in Theorem 4.3 that the spectrum is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$. The condition on the bilaplacian required as in [28, Theorem 2] follows automatically since the horospheres have constant mean curvature. In fact, along the way we prove that equality holds in Cheeger’s inequality [6, §4.3, Theorem 3] for asymptotically harmonic manifolds and that Cheeger’s constant is equal to $h$ in Proposition 4.1. Indeed, the proof shows that it is sufficient for horospheres to have constant mean curvature in one direction. Further, we prove in Theorem 4.7 that the spectrum is purely absolutely continuous for simply connected noncompact nonflat harmonic manifolds without focal points and is equal to $[h^{2}/4,\infty)$. These are harmonic manifolds of purely exponential volume growth [21, Theorems 32, 45]. Rank one symmetric spaces of noncompact type and other Damek-Ricci spaces are some examples of this class of manifolds. Since harmonic manifolds are asymptotically harmonic, that the spectrum is purely absolutely continuous and that it is contained in $[h^{2}/4,\infty)$ follow from Theorem 4.3. To prove that the spectrum is equal to $[h^{2}/4,\infty)$, we make use of the Fourier transform developed in this context in [3]. More precisely, using the unitarity of the spherical Fourier transform [3, Theorem 4.7], we can show in the usual way that the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$ is unitarily equivalent to the action of the Laplacian on radial functions to conclude that $[h^{2}/4,\infty)$ is contained in the spectrum.
Higher rank symmetric spaces of noncompact type are not asymptotically harmonic since the horospheres along each direction have constant mean curvature which depend on that direction. Nonetheless, we are able to obtain (Theorem 4.8) that the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\left\lVert\rho^{2}\right\rVert,\infty)$ by making use of the explicit expression for the mean curvature of horospheres and the strictly convex function that we construct in Theorem 3.1. To show that the spectrum is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$, we use the unitarity of just the spherical Fourier transform as in the case of harmonic manifolds. Note that in contrast to the case of harmonic manifolds where the unitarity of the Fourier transform is not known, it is classical that the Fourier transform in the case of symmetric spaces is unitary. As a consequence, the spectrum can also be completely determined for symmetric spaces via the unitary equivalence between the Laplacian and the multiplication operator $M_{(\left\lVert\lambda\right\rVert^{2}+\left\lVert\rho\right\rVert^{2})}$ given by the Fourier transform. It is worth remarking that in showing that the spectrum of the Laplacian on symmetric spaces of noncompact type is purely absolutely continuous, essential use is made only of the fact that they are manifolds without focal points and have negative Ricci curvature.
2. Preliminaries
Let $M$ be a complete Riemannian manifold and let $\gamma$ be a geodesic in $M$. A vector field $Y$ on $\gamma$ is called a Jacobi field if it satisfies the Jacobi equation
$Y^{{}^{\prime\prime}}+R(\gamma^{{}^{\prime}},Y)\gamma^{{}^{\prime}}=0.$
Here $\gamma^{{}^{\prime}}$ is the tangent vector field of $\gamma$, $R$ is the Riemann curvature tensor on $M$ and $Y^{{}^{\prime}}$ denotes the covariant derivative of $Y$ with respect to $\gamma^{{}^{\prime}}$ along $\gamma$. Two points $p$ and $q$ on $\gamma$ are said to be conjugate along $\gamma$ if there exists a nonzero Jacobi field which vanishes both at $p$ and at $q$. $M$ is called a manifold without conjugate points if for every geodesic $\gamma$ in $M$, no two points on $\gamma$ are conjugate along $\gamma$ or equivalently, if the exponential map is non-singular at every point in $M$. In fact, $M$ has no conjugate points if and only if $\exp_{p}:T_{p}M\to M$ is a covering map for each $p$ in $M$. So if $M$ is also simply connected, then $\exp_{p}$ is a diffeomorphism for all $p\in M$ and hence $M$ is necessarily noncompact. In this case, any two distinct points of $M$ are joined by a unique geodesic. A general reference for the basic Riemannian geometry used in this article is [7].
We now give a brief overview of Jacobi tensors and refer to [13] for a full account. Let $M$ be a complete simply connected Riemannian manifold without conjugate points and let ($SM$) $TM$ denote its (unit) tangent bundle. For $v\in SM$, let $\gamma_{v}$ denote the unit speed geodesic determined by $v$. A normal $(1,1)$-tensor $Y_{v}$ along the geodesic $\gamma_{v}$ is a bundle endomorphism of the normal bundle $N\gamma_{v}$ over $\gamma_{v}$ given by
$Y_{v}:{\mathbb{R}}\to\mathrm{End}N(\gamma_{v})=\bigcup\limits_{t\in{\mathbb{R}}}\mathrm{End}(N(\gamma_{v}(t))),\,N(\gamma_{v}(t)):=\left\{w\in T_{\gamma_{{}_{v}}(t)}M:{\langle w},{\gamma_{v}^{{}^{\prime}}(t)\rangle}_{\gamma_{{}_{v}}(t)}=0\right\}.$
A Jacobi tensor $Y_{v}$ is a normal $(1,1)$-tensor field along a geodesic $\gamma_{v}$ with transversal derivative (i.e. there exists some $t$ such that $ker~{}Y_{v}(t)\cap ker~{}Y_{v}^{{}^{\prime}}(t)=0$) satisfying the Jacobi equation
(2.1) $Y_{v}^{{}^{\prime\prime}}(t)+R(t)Y_{v}(t)=0$
Here $R(t)$ is the symmetric $(1,1)$-tensor along $\gamma_{v}$ induced by the curvature tensor $R$ and defined by
$R(t)w:=R(w,\gamma_{v}^{{}^{\prime}}(t))\gamma_{v}^{{}^{\prime}}(t)$
and $Y_{v}^{{}^{\prime}}$ is the covariant derivative of the tensor $Y_{v}$ along $\gamma_{v}$ with respect to the tangent vector field $\gamma_{v}^{{}^{\prime}}$. For each $s>0$ and $v\in SM$, let $D_{s,v}$ be a solution to (2.1) with the boundary conditions $D_{s,v}(0)=\mathrm{Id}$ and $D_{s,v}(s)=0$. Then $D_{v}:=\lim\limits_{s\to\infty}D_{s,v}$ exists and is referred to as the stable Jacobi tensor $D_{v}$ satisfying $D_{v}(0)=\mathrm{Id}$. A complete simply connected Riemannian manifold without conjugate points is said to be manifold of continuous asymptote if $v\to D_{v}^{{}^{\prime}}(0)$ is a continuous section of the vector bundle $\{(v,\phi):v\in SM,\phi\in\mathrm{End}(N(\gamma_{v})(0))\}$ over $SM$.
Given $t\in[0,\infty)$ and $v\in SM$, let $b_{t,v}:M\to\mathbb{R}$ denote the function defined by
$b_{t,v}(x):=d(x,\gamma_{v}(t))-t$
which, by triangle inequality, is increasing as a function of $t$. Since $b_{t,v}(\cdot)$ is absolutely bounded by $d(\gamma_{v}(0),\cdot)$, the function
$b_{v}:=\lim\limits_{t\to\infty}b_{t,v}$
is defined everywhere on $M$. The function $b_{v}$ is called the Busemann function associated to $v\in SM$. Recall that the level sets of Busemann functions are called horospheres. Horospheres are closed and hence are complete hypersurfaces with the induced metric. Recall that for Hadamard manifolds the Busemann functions are $C^{2}$ [12, §1.6.6]. Eschenburg [13] was able to relax the curvature restriction by making use of convergence properties of Jacobi tensors, and generalized the $C^{2}$-regularity of horospheres to manifolds of continuous asymptote. For such manifolds, [22, §11] gives a compactification known as the Busemann compactification. Let $o\in M$ be an arbitrary but fixed point. Then the Busemann boundary $\partial^{o}M$ is homeomorphic to $S_{p}M$ for any $p\in M$ [22, Proposition 11.4 (2)].
We see that for each $v\in SM$, the Hessian of the Busemann function $b_{v}$ and the stable Jacobi tensor $D_{v}$ are related as follows. The orthogonal projection $x^{\perp}(p,v)$ of $x\in T_{p}M$ onto $(\nabla b_{v}(p))^{\perp}$ will be denoted by $x^{\perp}$ for notational convenience.
Lemma 2.1.
For $p\in M$ and $x\in S_{p}M$, we have ${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp})$.
Proof.
As $\nabla b_{v}$ is a $C^{1}$ vector field with $\left\lVert\nabla b_{v}\right\rVert=1$, we have that $\nabla_{\nabla b_{v}}\nabla b_{v}=0$. Indeed ${\langle\nabla_{X}\nabla b_{v}},{Y\rangle}={\mathrm{H}}^{b_{v}}(X,Y)$ is symmetric in $X,Y$ and hence for every vector field $X$
${\langle\nabla_{\nabla b_{v}}\nabla b_{v}},{X\rangle}$ $={\langle\nabla_{X}\nabla b_{v}},{\nabla b_{v}\rangle}$
$=\frac{1}{2}X{\langle\nabla b_{v}},{\nabla b_{v}\rangle}=0$
In particular, we see that every integral curve of $\nabla b_{v}$ is a geodesic and hence complete. Then, the orthogonal decomposition $T_{p}M={\mathbb{R}}\nabla b_{v}\oplus(\nabla b_{v})^{\perp}$ enables us to write $x={\langle x},{\nabla b_{v}(p)\rangle}_{T_{p}M}\nabla b_{v}(p)+x^{\perp}$. Using $\nabla_{\nabla b_{v}}\nabla b_{v}=0$ and the bilinearity of ${\mathrm{H}}^{b_{v}}$ we have
${\mathrm{H}}^{b_{v}}(p)(x,x)={\mathrm{H}}^{b_{v}}(p)(x^{\perp},x^{\perp}).$
Notice in the above computation that when $X=\nabla b_{v}$ we have ${\mathrm{H}}^{b_{v}}(\nabla b_{v},\nabla b_{v})=0$ and hence $b_{v}$ is not strictly convex for any $v\in SM$. In fact, to the best of our knowledge, it is not known if Busemann functions are even convex for an arbitrary manifold of continuous asymptote. Using Lemma 2.1 and [13, Proposition 2, Theorem 1] we have the following relation between the stable Jacobi tensors and the Busemann functions in the case of manifolds of continuous asymptote: for $v\in SM$, $p\in M$, and $x\in T_{p}M$,
(2.2) ${\mathrm{H}}^{b_{v}}(p)(x,x)=-{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)x^{\perp}},{D_{\nabla b_{v}(p)}(0)x^{\perp}\rangle}$
Therefore for asymptotically harmonic manifolds, by taking traces in (2.2), we have $\Delta b_{v}=h=trace(D_{v}^{{}^{\prime}}(0))$ for every $v\in SM$.
3. A strictly convex function on a manifold without focal points
A complete simply connected Riemannian manifold $M$ is said to have no focal points if for every nontrivial Jacobi field $Y$ satisfying $Y(0)=0$, the map $t\mapsto\left\lVert Y(t)\right\rVert$ is strictly increasing for $t>0$ [25, Proposition 4]. In particular, any non-trivial Jacobi field cannot vanish more than once and hence every manifold without focal points is also a manifold without conjugate points. Note that $t\mapsto\left\lVert D_{v}(t)\right\rVert$ is monotonically decreasing for each $v\in SM$. This in particular shows that $M$ is of continuous asymptote [13, §5].
Henceforth in this section we assume that $M$ is a $n$-dimensional complete simply connected Riemannian manifold without focal points. To begin with let us recall the geodesic compactifications of $M$ and its connections to the Riemannian exponential map and the Busemann functions. We refer to [16] for more details on what follows. Given $v\in SM$ there exists a unique geodesic $\gamma_{v}$ in $M$ satisfying $\gamma_{v}^{{}^{\prime}}(0)=v$. Two vectors $v,w\in SM$ are said to be asymptotic if the distance between $\gamma_{v}(t)$ and $\gamma_{w}(t)$ is bounded for all $t\geq 0$. Let $M(\infty)$ be the set of all classes of asymptotic vectors and let $\overline{M}=M\cup M(\infty)$. $\overline{M}$ has a canonical topology with the following property: for any $p\in M$, the exponential map ${\mathrm{Exp}}_{p}:T_{p}M\to M$ extends uniquely to a homeomorphism $\Phi_{p}:\overline{T_{p}M}\to\overline{M}$ where $\overline{T_{p}M}\backslash T_{p}M$ is homeomorphic with the unit sphere $S_{p}M$ in $T_{p}M$. Then for any $p\in M$, there exists a unique geodesic $\gamma_{{}_{\nabla b_{v}(p)}}$ starting at $p$ asymptotic to $\gamma_{v}$. From our discussions in §2, we see that the Busemann boundary $\partial^{o}M$ and the geodesic compactification $M(\infty)$ are homeomorphic.
Further, for each $v\in SM$, the Busemann function $b_{v}$ is convex [13, Theorem 2] i.e., ${\mathrm{H}}^{b_{v}}$ is nonnegative definite for every $v\in SM$. Hence each horosphere is a convex submanifold in $M$. As remarked in the introduction, motivated by the construction on symmetric spaces of noncompact type of a strictly convex function, indicated in [10, penultimate paragraph p. 657], we now construct a strictly convex function out of the Busemann functions, using the geometry at infinity for manifolds without focal points discussed above.
Theorem 3.1.
Let $M$ be a complete non-compact simply connected Riemannian manifold with no focal points and negative Ricci curvature. Then $M$ admits a $C^{2}$-strictly convex function $F$ whose gradient is bounded.
Proof.
Let $o$ be an arbitrary but fixed point in $M$. Let $\{v_{k}\}$ be a countable dense subset in $S_{o}M$. Consider the Dirac measures $\delta_{v_{k}}$ associated to each $v_{k}$. Then ${\eta=\sum\limits_{k=1}^{\infty}\frac{1}{2^{k}}\delta_{v_{k}}}$ defines a probability measure on $S_{o}M$. We identify $S_{o}M$, $M(\infty)$ and $\partial^{o}M$ via the respective homeomorphisms. Let $F(p)=\displaystyle{\int\limits_{M(\infty)}b_{v}(p)~{}\mathrm{d}\eta(v)}$. By dominated convergence theorem, the gradient of $F$ at $p$ is given by $\nabla F(p)=\displaystyle{\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)}$. Clearly,
$\left\lVert\nabla F(p)\right\rVert=\left\lVert\int\limits_{M(\infty)}\nabla b_{v}(p)~{}\mathrm{d}\eta(v)\right\rVert\leq\int\limits_{M(\infty)}\left\lVert\nabla b_{v}(p)\right\rVert~{}\mathrm{d}\eta(v)=1,\quad\text{as}~{}\left\lVert\nabla b_{v}\right\rVert=1$
i.e., $F$ has bounded gradient. In particular, $F$ is Lipschitz with Lipschitz constant $1$. For $p\in M$ and $w\in S_{p}M$, using (2.2) and that ${\mathrm{H}}^{b_{v}}$ is nonnegative definite we have
$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)$ $=\left|\int\limits_{M(\infty)}{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}~{}\mathrm{d}\eta(v)\right|$
$\leq\int\limits_{M(\infty)}|{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}|~{}\mathrm{d}\eta(v)$
[…]
$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert~{}\mathrm{d}\eta(v)\quad(D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)~{}\text{is a bounded linear map for}~{}v\in S_{p}M).$
The restriction of the continuous map $w\mapsto\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert$ on $SM$ to $S_{p}M$ shows that there exists some $L>0$ such that $\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $w\in S_{p}M$. In particular, $\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $v\in S_{o}M$. So another application of the dominated convergence theorem shows that
${\mathrm{H}}^{F}(p)(w,w)=\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v).$
The convexity of $F$ follows from the convexity of the Busemann functions. We now claim that $F$ is, in fact, strictly convex. Suppose there exists $p\in M$ and $0\neq w\in T_{p}M$ such that ${\mathrm{H}}^{F}(p)(w,w)=0$ then
$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)=0$
Since $p\mapsto{\mathrm{H}}^{b_{v}}(p)(w,w)$ is a nonnegative continuous function in $v$, we have ${\mathrm{H}}^{b_{v}}(p)(w,w)=0$ for all $v\in S_{o}M$. Using (2.2),
(3.1) ${\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}},{D_{\nabla b_{v}(p)}(0)w^{\perp}\rangle}=0,\quad\text{for all}~{}v\in S_{o}M.$
It is known from [13, Section 5] that for $v\in SM$ and all parallel vector fields $x$ normal to $\gamma_{v}$, $\frac{d}{dt}(\left\lVert D_{v}(t)x\right\rVert^{2})\leq 0$ i.e., the function $f_{v,x}(t)=\frac{1}{2}\left\lVert D_{v}(t)x\right\rVert^{2}$ is $C^{2}$ and monotonically decreasing. From (3.1) we see that for each $v\in S_{o}M$, $f_{\nabla b_{v}(p),w^{\perp}}$ has a critical point at 00. We claim that $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. If not, then 00 is either a local minimum or local maximum. Suppose 00 is a local minimum then there exists a $\delta$-neighbourhood of 00 such that $f_{\nabla b_{v}(p),w^{\perp}}(0)\leq f_{\nabla b_{v}(p),w^{\perp}}(t)$ for all $t\in(-\delta,\delta)$. Since $f$ is monotonically decreasing we have $f_{\nabla b_{v}(p),w^{\perp}}(t)=f_{\nabla b_{v}(p),w^{\perp}}(0)$ for all $t\in(0,\delta)$. Hence $f_{\nabla b_{v}(p),w^{\perp}}$ is constant on $(0,\delta)$. Observe that $f_{\nabla b_{v}(p),w^{\perp}}$ is $C^{2}$ and so $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=\lim\limits_{t\to 0_{+}}f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(t)=0$, a contradiction. A similar argument could be carried out to show that 00 cannot be a local maximum of $f_{\nabla b_{v}(p),w^{\perp}}$. Hence 00 is a degenerate critical point of $f_{\nabla b_{v}(p),w^{\perp}}$ i.e, $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. Then the Jacobi equation for $D_{\nabla b_{v}(p)}$,
$D_{\nabla b_{v}(p)}^{{}^{\prime\prime}}(0)+R_{\nabla b_{v}(p)}(0)D_{\nabla b_{v}(p)}(0)=0$
gives
00 0 $=f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)$
$={\langle D_{\nabla b_{v}(p)}^{{}^{\prime\prime}}(0)w^{\perp}},{w^{\perp}\rangle}+\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
$=-{\langle{R_{\nabla b_{v}(p)}}(0)w^{\perp}},{w^{\perp}\rangle}+\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
$\implies{\langle R_{\nabla b_{v}(p)}(0)w^{\perp}},{w^{\perp}\rangle}$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
(3.2) $\implies K(\nabla b_{v}(p),w^{\perp})$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}\geq 0,\quad\text{for all $v\in S_{o}M$}$
where $K$ denotes the sectional curvature corresponding to the plane spanned by $\nabla b_{v}(p)$ and $w^{\perp}$. Let $x\in S_{p}M$ be such that $x\perp w$. Since $S_{p}M$ is homeomorphic to $M(\infty)$ there exists a unique $v_{o}\in S_{o}M$ such that $\gamma_{v_{o}}$ and $\gamma_{x}$ are asymptotic. As $\{v_{k}\}$ is dense in $S_{o}M$ there exists a subsequence $(v_{k_{m}})$ in $S_{o}M$ converging to $v_{o}$. For each $m$ we have $\nabla b_{v_{{}_{k_{m}}}}(p)\in S_{p}M$ such that $\gamma_{\nabla b_{v_{{}_{k_{m}}}}(p)}$ and $\gamma_{v_{{}_{k_{m}}}}$ are asymptotic and so $v_{{}_{k_{m}}}$ and $\nabla b_{v_{{}_{k_{m}}}}(p)$ correspond to the same element in $M(\infty)$. It then follows that the sequence $(\nabla b_{v_{{}_{k_{m}}}}(p))$ converges to $x$ in $S_{p}M$. We denote the orthogonal projection of $w$ onto $(\nabla b_{v_{{}_{k_{m}}}}(p))^{\perp}$ by $w^{\perp}(m)$. Then
$\left\lVert w-w^{\perp}(m)\right\rVert^{2}$ $=\left\lVert{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}\nabla b_{v_{k_{m}}}(p)\right\rVert^{2}$
$=|{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}|^{2}\left\lVert\nabla b_{v_{k_{m}}}(p)\right\rVert^{2}$
$=|{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}|^{2}$
$\xrightarrow{m\to\infty}0\quad\text{as $\nabla b_{v_{k_{m}}}(p)\to x$ and $x\perp w$}.$
Thus the sequence $(w^{\perp}(m))$ converges to $w$ as $m\to\infty$ and from (3) we have,
$K(\nabla b_{v_{k_{m}}}(p),w^{\perp}(m))\to K(x,w)\geq 0,\quad\text{as $m\to\infty$}$
Now for $\{x_{1},\cdots,x_{n-1}\}$ orthonormal vectors in $T_{p}M$ such that $x_{i}\perp w$, $Ricc(w,w)=\sum\limits_{i=1}^{n}K(x_{i},w)\geq 0$. This is a contradiction to the hypothesis that $M$ has negative Ricci curvature. Hence ${\mathrm{H}}^{F}(p)(w,w)>0$ for every $p\in M$ and every $0\neq w\in T_{p}M$ proving the strict convexity of $F$. ∎
While there are several geometrically motivated convex functions on a Riemannian manifold, typically they have low regularity. Even in situations where such convex functions have higher regularity they often do not have other useful properties such as having bounded gradient. On the one hand, distance functions and Busemann functions are examples of convex functions with bounded gradient having low regularity. On the other hand, on complete simply connected non-compact Riemannian manifolds without focal points, the squared distance function is a smooth strictly convex function on $M$ [11, Lemma 1] whose gradient is not bounded. It is in this context that we have the following approximation result for strictly convex functions with bounded gradient. Note that the smoothing approximation fails if we drop the assumption of strict convexity [26]. Recall that a $C^{2}$-function is strictly convex if its Hessian is positive definite. It was proved in [17, Theorem 2(a)] that a Lipschitz $C^{2}$-function which is strictly convex admits a smooth strictly convex approximation which is also Lipschitz. From this and using the fact that Lipschitz functions have bounded gradient we have the following corollary to our Theorem 3.1.
Corollary 3.2.
Let $M$ be as in Theorem 3.1. Then for any $C>1$, $M$ admits a smooth strictly convex function $f$ such that $\left\lVert\nabla f\right\rVert\leq C$.
The property of having bounded gradient is useful in applications as can be seen below in Corollary 3.3 and Theorem 4.7.
Let $(M,g,J)$ be a Kähler manifold as in the introduction. The complex analogue of convex functions is given by the notion of plurisubharmonic functions. These are defined using the complex Hessian called the Levi form. In local coordinates, the Levi form of a function $f$ is given by
$Lf=4\sum\limits_{i,j=1}^{n}\frac{\partial^{2}f}{\partial z_{i}\partial\overline{z}_{j}}\mathrm{d}z_{i}\otimes\mathrm{d}\overline{z}_{j}$
where $(z_{1},\cdots,z_{n})$ is a local holomorphic coordinate system. Let $(x_{1},y_{1},\cdots,x_{n},y_{n})$ be the corresponding normal coordinates with $z_{k}=x_{k}+iy_{k}$, $k=1,\cdots,n$. A real-valued function on a Kähler manifold is said to be (strictly) plurisubharmonic, if the Levi form is positive (definite) semi-definite. Since $M$ is Kähler, the Levi form $L_{f}$ of $f$ is expressed in terms of Hessian of $f$ using the real coordinates as
$L_{f}={\mathrm{H}}^{f}\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{1}}\right)+{\mathrm{H}}^{f}\left(\frac{\partial}{\partial y_{1}},\frac{\partial}{\partial y_{1}}\right).$
Therefore if $f$ is a strictly convex $C^{2}$-function, then $f$ is strictly plurisubharmonic. As an application to Corollary 3.2, using [24, Lemma 4.2 (i)], we show the following volume growth comparison for metric balls.
Corollary 3.3.
If $M$ as in Theorem 3.1 is also Kähler of (complex) dimension $m$, then $M$ admits a smooth strictly plurisubharmonic function with bounded gradient. Moreover, for each $p\in M$ there exists a constant $C_{2}(p)>0$ such that ,
$V_{p}(r)\geq C_{2}(p)r^{m},\quad r\geq 1$
where $V_{p}(r)$ is the volume of a metric ball of radius $r$ about $p$ in $M$.
4. Spectral theoretic applications
We begin with a sketch of the basic functional analytic setup concerning the spectrum of the Laplacian following [4]. Let $M$ be a complete Riemannian manifold. Then the (positive) Laplacian $\Delta$ is an essentially self-adjoint operator with the space of compactly supported smooth functions on $M$ as the domain. We also denote its unique self-adjoint extension, having the domain $\mathcal{D}(\Delta):=\{f\in L^{2}(M):\Delta f\in L^{2}(M)\}$, by $\Delta$. Using the functional calculus for $\Delta$ we have the family of spectral projection operators $\prod_{E}:=\chi_{E}(\Delta)$ for each Borel subset $E$ of ${\mathbb{R}}$ referred to as the resolution of $\Delta$. The support of these projections and the spectrum of $\Delta$ are related in the following way. We say that $\lambda\in{\mathbb{C}}$ is in the spectrum $\sigma(\Delta)$ of $\Delta$ if and only if $\prod_{(\lambda-\epsilon,\lambda+\epsilon)}\neq 0$ for all $\epsilon>0$. If $\prod_{\{\lambda\}}\neq 0$, then $\lambda$ is an eigenvalue of $\Delta$ and the range of $\prod_{\{\lambda\}}$ is the corresponding eigenspace. The spectral theorem gives a unitary equivalence between $\Delta$ and a multiplication operator on a suitable Hilbert space of the form $\mathrm{L}^{2}(X,\mu)$, where $\mu$ is a $\sigma-$ finite Borel measure on $X$. By the Lebesgue decomposition, $\mu$ decomposes as $\mu=\mu_{pp}+\mu_{ac}+\mu_{sc}$ into pure point measure, absolutely continuous and singularly continuous measures with respect to the Lebesgue measure, and consequently we have that
(4.1) $\mathrm{L}^{2}(X,\mu)=\mathrm{L}^{2}(X,\mu_{pp})\oplus\mathrm{L}^{2}(X,\mu_{ac})\oplus\mathrm{L}^{2}(X,\mu_{sc}).$
This induces a decomposition $\mathrm{L}^{2}(M)=\mathcal{H}_{pp}\oplus\mathcal{H}_{ac}\oplus\mathcal{H}_{sc}$ from which we have the associated decomposition of the spectrum of $\Delta$ namely, $\sigma(\Delta)=\overline{\sigma_{pt}(\Delta)}\cup\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ where
$\sigma_{ac}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{ac}}}),\quad\sigma_{sc}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{sc}}}),\quad\overline{\sigma_{pt}(\Delta)}=\sigma(\Delta_{|_{\mathcal{H}_{pp}}})$
Here we take the closure for the set of all eigenvalues as it is not necessarily closed. We refer to $\sigma_{ac}(\Delta)$ and $\sigma_{sc}(\Delta)$ as the absolutely continuous and singularly continuous parts of $\sigma(\Delta)$ respectively and their union as the continuous part $\sigma_{cont}(\Delta)$ of $\sigma(\Delta)$.
In order to analyse the nature of the spectrum of the Laplacian, we invoke [28, Theorem 2]. This result demands boundedness of the gradient and of the Laplacian of a strictly convex function as well as the additional constraint of nonpositivity on the bilaplacian and in turn yields the absolute continuity of the spectrum. Consequently, it enables us to exclude the existence of the point spectrum and the singularly continuous spectrum. As pointed out in [28], the seemingly unnatural additional condition that $\Delta^{2}F\leq 0$ is, in fact, necessary. Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed there. The proof of absolute continuity of the spectrum is then derived from an estimate implementing Kato’s theory of $H-$smooth operators in Hilbert spaces. We shall observe below that the condition on the bilaplacian is satisfied for $M$ as in Theorem 3.1 which are in addition asymptotically harmonic.
We first briefly discuss some facts regarding the bottom $\lambda_{0}(M)$ of the spectrum for asymptotically harmonic manifolds. For a complete noncompact Riemannian manifold $M$ it is characterized by [6, §4.4]
(4.2) $\lambda_{0}(M)=\inf\limits_{f\neq 0}\frac{\int\limits_{M}|\nabla f|^{2}}{\int\limits_{M}|f|^{2}}$
where $f$ varies over the completion $\mathfrak{H}(M)$ of $C_{c}^{\infty}(M)$ with respect to the norm $\left\lVert f\right\rVert=\left\lVert f\right\rVert_{\rm{L}^{2}(M)}+\left\lVert\nabla f\right\rVert_{\rm{L}^{2}(M)}$. Let $|\Omega|$ denote the volume of a submanifold $\Omega$ in $M$. Then the Cheeger’s constant of $M$ is given by [6, §4.3]
$h_{Cheeger}(M):=\inf\limits_{\Omega\subset M}\frac{|\partial\Omega|}{|\Omega|}$
where the infimum varies over all open, connected submanifolds $\Omega$ of $M$ with compact closure and smooth boundary $\partial\Omega$. The importance of this global geometric invariant is illustrated by the following fundamental inequality relating it to the global analytic invariant $\lambda_{0}(M)$ [6, §4.3, Theorem 3],
(4.3) $\frac{h_{Cheeger}^{2}(M)}{4}\leq\lambda_{0}(M).$
Proposition 4.1.
Let $M$ be an asymptotically harmonic manifold with horospheres of mean curvature $h$ i.e. $\Delta b_{v}=h$ for all $v\in SM$. Then we have $h_{Cheeger}=h$ and $\lambda_{0}(M)=h^{2}/4$.
Proof.
For an open connected $\Omega$ in $M$ with compact closure and smooth boundary $\partial\Omega$ we have, by Stokes’ theorem, that
$h|\Omega|=\int\limits_{\Omega}\Delta b_{v}=\left|\int\limits_{\partial\Omega}{\langle\nabla b_{v}},{\nu\rangle}\right|\leq|\partial\Omega|,\quad(\text{since$\left\lVert\nabla b_{v}\right\rVert=1$})$
where $\nu$ is the outer unit normal vector field of $\Omega$ along $\partial\Omega$. Thus,
(4.4) $h\leq\inf\limits_{\Omega\subseteq M}\frac{|\partial\Omega|}{|\Omega|}=h_{Cheeger}$
where the infimum varies over $\Omega$ as above in $M$. Let $o\in M$ be arbitrary. Since $M$ is simply connected without conjugate points the exponential at $o$ is a diffeomorphism. Let $d(o,\cdot):M\to{\mathbb{R}}$ denote the distance from $o$. By taking the sequence of functions $f_{n}(x)=e^{-\frac{1}{2}(h+\frac{1}{n})d(o,x)}$ in $\mathfrak{H}(M)$, $n\in{\mathbb{N}}$, we see from (4.2) that
(4.5) $\lambda_{0}(M)\leq\frac{h^{2}}{4}$
Combining (4.3), (4.4), (4.5) we have
(4.6) $h=h_{Cheeger}\quad\text{and}\quad\lambda_{0}(M)=\frac{h^{2}}{4}.$
We remark here that for the proposition above to hold it suffices that there is a $v\in SM$ such that the corresponding horospheres have constant mean curvature.
Lemma 4.2.
Let $M$ and $F$ be as in Theorem 3.1. Further, assume that $M$ is asymptotically harmonic with $h$ as the constant mean curvature of the horospheres. Then $F$ has bounded Laplacian and $\Delta^{2}F=0$.
Proof.
Since $\Delta$ is the trace of Hessian, by choosing an orthonormal basis $\{E_{i}(p)\}_{i=1}^{n}$ of $T_{p}M$ we have
$\sum\limits_{i=1}^{n}{\mathrm{H}}^{F}(p)(E_{i}(p),E_{i}(p))$ $=\sum\limits_{i=1}^{n}\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
$=\int\limits_{M(\infty)}\sum\limits_{i=1}^{n}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
$\Delta F(p)$ $=\int\limits_{M(\infty)}\Delta b_{v}(p)~{}\mathrm{d}\eta(v)=h$
and the lemma follows. ∎
Theorem 4.3.
If $M$ as in Theorem 3.1 is also asymptotically harmonic with $h$ as the constant mean curvature of the horospheres, then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\frac{h^{2}}{4},\infty)$.
Proof.
The absolute continuity of the spectrum follows by combining [28, Theorem 2] with Lemma 4.2. That the spectrum of Laplacian is contained in $[\frac{h^{2}}{4},\infty)$ follows from (4.6). ∎
Next we briefly describe harmonic manifolds, only mentioning aspects which are needed in the sequel and to set notation, and we refer to [1, Chapter $6$], and to the comprehensive exposition [21] for the details. A function $f$ on a Riemannian manifold $(M,g)$ is said to be radial around a point $o$ if $f$ is constant on geodesic spheres $S_{o}(r)$ of radius $r$ centered at $o$. $M$ is said to be harmonic if the volume density in normal coordinates centered at $o$ is a radial function. Given a continuous function $f$, its radialisation about $o$ given by
$R_{o}(f)(z):=\int\limits_{S_{o}(r)}f(y){\rm{d}}\sigma^{r}(y)$
where $\sigma^{r}$ denotes surface area measure on $S_{o}(r)$ (induced from the Riemannian metric on $M$), normalized to have mass one. Note that for $f$ radial around $o$, we have that $R_{o}f=f$. Harmonicity of the manifold can be seen to be equivalent to the fact that the radialisation operator $R_{o}$ commutes with the Laplacian $\Delta$. Harmonic manifolds are necessarily Einstein manifolds with Ricci curvature $Ric(g)=\kappa g$. Recall that we have the following description of simply connected harmonic manifolds such that $\kappa\geq 0$:
  • (a)
    If $\kappa>0$, then $M$ is a compact rank one symmetric space
  • (b)
    If $\kappa=0$, then $M$ is Euclidean.
Thus simply connected noncompact nonflat harmonic manifolds necessarily have constant negative Ricci curvature. For every $v\in SM$, the Busemann function $b_{v}$ satisfies $\Delta b_{v}=h$ (see for instance [5, §2, p. 33]) and so it follows that every harmonic manifold is also asymptotically harmonic. Consequently we obtain the following corollary to Theorem 4.3.
Corollary 4.4.
Let $M$ be a complete simply connected noncompact nonflat harmonic manifold $M$ without focal points. Then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$.
We now proceed to show that the spectrum is actually equal to the interval $[h^{2}/4,\infty)$, by making use of the unitarity of the spherical Fourier transform for radial functions on harmonic manifolds of purely exponential volume growth studied in [3]. Notice that in [3] unitarity is shown only in dimensions greater than $5$. However, the only simply connected noncompact nonflat harmonic manifolds of dimension less than equal to $5$ are Riemannian symmetric spaces of noncompact type. This case is handled in Theorem 4.8 below. First, we observe from [21, Theorems 32, 45] that, noncompact harmonic manifolds without focal points fall into the class of noncompact harmonic manifolds of purely exponential volume growth. Let $o$ be an arbitrary but fixed point in $M$. Then for $f\in C_{c}^{\infty}(M)$ which is radial about $o$, the spherical Fourier transform ${\mathcal{F}}^{o}f$ is given by
${\mathcal{F}}^{o}(f)(\lambda,v)=\int\limits_{M}f(y)(\phi_{\lambda}\circ d_{o})(y){\rm{d}}y$
for $\lambda\in{\mathbb{C}}$ and $v\in S_{o}M$. Here $\phi_{\lambda}$ is the unique function on $[0,\infty)$ with $\phi_{\lambda}(0)=1$ and satisfying $\Delta(\phi_{\lambda}\circ d_{o})=(\lambda^{2}+\frac{h^{2}}{4})(\phi_{\lambda}\circ d_{o})$, where $d_{o}$ denotes the distance from $o$. Let ${\rm{L}}_{o}^{2}(M)$ denote the closed subspace of ${\rm{L}}^{2}(M)$ consisting of those functions which are radial around $o$ in ${\rm{L}}^{2}(M)$. It is known from [3, Theorem 4.7] that the ${\mathcal{F}}^{o}$ is an unitary operator from ${\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$. Here the $c$-function coincides with the Harish Chandra $c$-function on $Im(\lambda)\leq 0$, for rank one symmetric spaces of noncompact type.
We recall that for a real-valued measurable function $f$ on a $\sigma-$finite measure space $(X,\mu)$, the multiplication operator $M_{f}$ on ${\rm{L}}^{2}(X,\mu)$ is self-adjoint with domain $\mathcal{D}(M_{f}):=\{g\in{\rm{L}}^{2}(X,\mu):M_{f}g\in{\rm{L}}^{2}(X,\mu)\}$. Let $m$ be the function on $[0,\infty)$ defined by $m(\lambda)=\lambda^{2}+\frac{h^{2}}{4}$. Then we see that the multiplication operator $M_{m}$ on ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$ is self-adjoint on
$\mathcal{D}(M_{m})=\{f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda):M_{m}f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)\}.$
Moreover, the spectrum of a multiplication operator $M_{m}$ is equal to the essential range of $m$ [4, Theorem 4.5]. Here
$ess-range(m):=\{x\in{\mathbb{R}}:\mu(m^{-1}(B(x,\epsilon)))>0~{}\forall~{}\epsilon>0\}.$
Now to determine the spectrum of $M_{m}$ we argue as follows. For any $x\in{\mathbb{R}}$ and $\epsilon>0$,
$m^{-1}(B(x,\epsilon))=\{\lambda\in[0,\infty):x-\epsilon<\lambda^{2}+\frac{h^{2}}{4}<x+\epsilon\}$
Note that when $x<\frac{h^{2}}{4}$ and $\epsilon=-(x-\frac{h^{2}}{4})$, $m^{-1}(B(x,\epsilon))$ is empty and so $x\notin ess-range(m)$. Equivalently, $ess-range(m)\subseteq[\frac{h^{2}}{4},\infty)$. Now for any $x\in[\frac{h^{2}}{4},\infty)$ and $\epsilon>0$ we have,
$m^{-1}(B(x,\epsilon))=\begin{cases}\left[0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon<0\\ \\ \left[\sqrt{x-\frac{h^{2}}{4}-\epsilon},\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)&\text{if}~{}x-\frac{h^{2}}{4}-\epsilon\geq 0.\\ \end{cases}$
Using the estimates on $|c(\lambda)|^{-1}$ given in [3, §4.1] we see that, for $C>0$ and $K=\sqrt{x-\frac{h^{2}}{4}+\epsilon}>0$,
$\int\limits_{m^{-1}(B(x,\epsilon))}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\frac{C_{0}}{C}\int\limits_{m^{-1}(B(x,\epsilon))}|\lambda|^{2}{\rm{d}}\lambda>0.$
It follows that $ess-range(m)=[\frac{h^{2}}{4},\infty)$ and so the spectrum of $M_{m}$ is equal to $[\frac{h^{2}}{4},\infty)$. For functions that are not necessarily radial, a notion of Fourier transform, relative to a point, was introduced in [3, §5] where it was shown to be an isometry but the unitarity is not known. This Fourier transform coincides with the spherical Fourier transform ${\mathcal{F}}^{o}$ for radial functions [3, Proposition 5.3]. Then by using [5, Lemma 6.5], we see that
${\mathcal{F}}^{o}\Delta f=M_{m}{\mathcal{F}}^{o}f,\quad f\in{\rm{L}}_{o}^{2}(M).$
Since ${\mathcal{F}}^{o}$ is unitary and $\mathcal{D}(M_{m})={\mathcal{F}}^{o}(\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M))$, we have the following lemma.
Lemma 4.5.
The spectrum of the Laplacian $\Delta$ on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ is equal to $[h^{2}/4,\infty)$.
We now claim that,
Lemma 4.6.
$[h^{2}/4,\infty)\subseteq\sigma(\Delta)$ where $\sigma(\Delta)$ is the spectrum of $\Delta$ with domain $\mathcal{D}(\Delta)$.
Proof.
If $\mu$ is not in $\sigma(\Delta)$, then $\Delta-\mu:\mathcal{D}(\Delta)\to{\rm{L}}^{2}(M)$ is invertible with bounded inverse. In particular, $\Delta-\mu$ is injective on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. If $f\in{\rm{L}}_{o}^{2}(M)$, then there exists a unique $g$ in $\mathcal{D}(\Delta)$ such that $(\Delta-\mu)g=f$. We have
$(\Delta-\mu)g=f=R_{o}f=R_{o}(\Delta-\mu)g=(\Delta-\mu)R_{o}g$
and by uniqueness $R_{o}g=g$ i.e., $\Delta-\mu$ is surjective from $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}_{o}^{2}(M)$ and hence invertible. Since $(\Delta-\mu)^{-1}$ is bounded on ${\rm{L}}^{2}(M)$ and ${\rm{L}}_{o}^{2}(M)$ is a closed subspace, $(\Delta-\mu)^{-1}$ is also bounded on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. The claim now follows from Lemma 4.5. ∎
The equality of the spectrum now follows from Corollary 4.4 and Lemma 4.6. Therefore, we have the following theorem.
Theorem 4.7.
Let $M$ be a simply connected noncompact nonflat harmonic manifold without focal points. Then the spectrum of Laplacian is purely absolutely continuous and is equal to $[h^{2}/4,\infty)$.
[…]
$R_{v}e_{i}=-R(v,e_{i})v=\lambda_{i}(v)e_{i}.$
From the structure theory for noncompact connected real semisimple Lie algebras, we have the set of all positive roots denoted by $\sum^{+}$. Let $m_{\alpha}$ denote the multiplicity of $\alpha\in\sum^{+}$. The finite reflection group $W$ generated by the roots is known as the Weyl group. We denote the tangent vector that is dual to $2\rho=\sum\limits_{\alpha\in\sum^{+}}m_{\alpha}\alpha$ also by $2\rho$. The arguments in [27, Lemma 1] show that for each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function given by $\sum\limits_{i}\sqrt{\lambda_{i}(v)}={\langle v},{2\rho\rangle}$. This constant is the mean curvature of horospheres with respect to $b_{v}$. Here we argue as in Proposition 4.1 by using Stoke’s theorem to get a lower bound for Cheeger’s constant namely, ${\langle v},{2\rho\rangle}\leq h_{Cheeger}$ for all $v\in S_{o}M$. In particular, for $v=\frac{\rho}{\left\lVert\rho\right\rVert}$ we have $\left\lVert\rho\right\rVert\leq h_{Cheeger}$ and hence by Cheeger’s inequality we obtain
(4.7) $\left\lVert\rho\right\rVert^{2}\leq\lambda_{0}(M).$
Observe that $M$ has negative Ricci curvature [12, §2.14, p. 96] and has no focal points. With notations as in Theorem 3.1 we have,
  1. (i)
    $\left\lVert\nabla F\right\rVert\leq 1$ since $\left\lVert\nabla b_{v}\right\rVert=1$ for all $v\in S_{o}M$.
  2. (ii)
    From $\Delta b_{v}={\langle v},{2\rho\rangle}$ for all $v\in S_{o}M$, we have $|\Delta b_{v}|\leq\left\lVert 2\rho\right\rVert$. Thus $\Delta b_{v}$ is bounded independent of $v\in S_{o}M$ by $\left\lVert 2\rho\right\rVert$ and so $|\Delta F|\leq\left\lVert 2\rho\right\rVert$.
  3. (iii)
    For each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function ${\langle v},{2\rho\rangle}$ and so we have $\Delta^{2}b_{v}=0$. It follows that, $\Delta^{2}F=0$.
The absolute continuity of the spectrum $\sigma(\Delta)$ is then established by invoking [28, Theorem 2]. Further, using (4.7) we have,
Theorem 4.8.
Let $M$ be a symmetric space of noncompact type. Then the spectrum of the Laplacian on $M$ is absolutely continuous and is equal to $[\left\lVert\rho^{2}\right\rVert,\infty)$.
Here we use the spherical Fourier transform for the $K$-bi-invariant functions on $G$ to prove the equality of the spectrum. The radial part of the Laplacian $\Delta$ for the action of $K$ on $G/K$ is given by [19, §4.5, p. 427],
(4.8) $\Delta_{rad}\phi_{\lambda}=({\langle\lambda},{\lambda\rangle}+{\langle\rho},{\rho\rangle})\phi_{\lambda},\quad\lambda\in\mathfrak{a}_{{\mathbb{C}}}^{\ast}$
where $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$ the complexification of the dual $\mathfrak{a}^{\ast}$ of $\mathfrak{a}$ and $\phi_{\lambda}(gK)=\displaystyle{\int\limits_{K}e^{(i\lambda+\rho)A(kg)}~{}{\rm{d}}k}$ is the unique spherical function on $G$ with the normalization $\phi_{\lambda}(eK)=1$ and $A$ is referred to as the Iwasawa projection $A:G\to{\mathfrak{a}}$. Two spherical functions are identical up to the action by the Weyl group $W$ on $\mathfrak{a}_{{\mathbb{C}}}^{\ast}$. Let us denote the space of smooth compactly supported $K-$invariant functions on $G/K$ by $C_{c}^{\infty}(K\backslash G/K)$ and its closure in ${\rm{L}}^{2}(G/K)$ by ${\rm{L}}_{rad}^{2}(G/K)$. Using the inversion formula for the spherical transform [19, Chapter 4, Theorem 7.5] the map $f\mapsto\widehat{f}$
$\widehat{f}(\lambda)=\int\limits_{G/K}f(gK)\phi_{\lambda}(gK),\quad f\in C_{c}^{\infty}(K\backslash G/K)$
[…]
(4.9) $\int\limits_{m^{-1}(B(x,\epsilon))}|\textbf{c}(\lambda)|^{-2}{\rm{d}}\lambda>0.$
for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$. Note that $|\textbf{c}(\lambda)|^{2}$ defined on ${\mathfrak{a^{\ast}}}$ is invariant under the action of $W$ i.e., $|\textbf{c}(w\lambda)|^{2}=|\textbf{c}(\lambda)|^{2}$ for all $w\in W$ [19, §IV.7, Equation 16] and hence descends to a function on ${\mathfrak{a}^{\ast}}/W$. Since $\lambda\mapsto\left\lVert\lambda\right\rVert^{2}$ is continuous and
$m^{-1}(B(x,\epsilon))=\left\{\lambda\in{\mathfrak{a}^{\ast}}/W:x-\epsilon-\left\lVert\rho\right\rVert^{2}<\left\lVert\lambda\right\rVert^{2}<x+\epsilon-\left\lVert\rho^{2}\right\rVert\right\}$
we see that $m^{-1}(B(x,\epsilon))$ contains an open ball in $\mathfrak{a}^{\ast}/W$. Using [15, Chapter 3, p. 101, Theorem 6.3.4], we have that the map $\lambda\mapsto|\textbf{c}(\lambda)|^{-2}$ is analytic on $\mathfrak{a}^{\ast}$ and thereby analytic on $\mathfrak{a}^{\ast}/W$. Thus, (4.9) holds for $x\geq\left\lVert\rho\right\rVert^{2}$ and $\epsilon>0$. This, in particular, shows that the spectrum of $\Delta_{rad}$ is equal to $[\left\lVert\rho\right\rVert^{2},\infty)$. We argue as in Lemma 4.6 to show the equality in Theorem 4.8.
5. Funding
This work was supported by the Centre for Operator Algebras, Geometry, Matter and Spacetime, Ministry of Education, Government of India through Indian Institute of Technology Madras [SB22231267MAETWO008573 to A.P. and S.B.]; and by the National Board for Higher Mathematics, Department of Atomic Energy, Government of India [Ref.No:0203/11/2019-R&D-II to S.B.]