### v2

# Convexity on manifolds without focal points and applications

Abstract.
In this article, we study strict convexity on simply connected Riemannian manifolds without focal points, a class of manifolds which contains not only Hadamard manifolds but also examples of manifolds whose sectional curvatures change sign. Using the geometrically defined convex functions on such manifolds, namely the distance function and the Busemann function, strictly convex functions with various desired additional properties are constructed. From the former, we derive interesting consequences such as the continuity of the isoperimetric profile function without conditions on the sectional curvature and, if the manifolds are also Kähler, Steinness as well as a lower bound on the volume growth of metric balls. Our primary applications concern the spectrum of the Laplacian. We prove that the absolutely continuous part of the spectrum contains a certain infinite interval assuming only the existence of a point with respect to which the radial curvatures are nonpositive. This is a substantial generalisation of the corresponding result for Hadamard manifolds. In the latter case of Busemann functions, we use the geometry at infinity to give a new construction of a strictly convex function assuming that the Ricci curvature is negative. We then apply this to show that the spectrum is purely absolutely continuous on a class of manifolds for which all the horospheres have the same constant mean curvature. In particular, for a well-studied subclass for which the condition on Ricci curvature is automatically satisfied, we completely determine the spectrum. This construction also enables us to recover the classical result on the spectrum in the case of symmetric spaces of noncompact type and higher rank even though the mean curvatures of horospheres there are not constant.
1. Introduction
Convex functions have long played a fundamental role in Riemannian geometry. By the convexity of a continuous real-valued function on a Riemannian manifold we mean that it is convex in the usual sense along any geodesic. The mere existence or nonexistence of convex functions have geometric and topological implications. Nonconstant convex functions on complete Riemannian manifolds exist only on those with infinite volume. In particular, they don’t exist on compact Riemannian manifolds. There are profound structural results arising from the interaction of curvature and convexity. Under appropriate curvature conditions there are natural geometrically defined convex functions on a Riemannian manifold. On simply connected Riemannian manifolds of nonpositive sectional curvature, called Hadamard manifolds, the square of the distance function from a fixed but arbitrary point and the Busemann functions are examples of convex functions which play a crucial role in the geometry of these manifolds. While the former is a smooth strictly convex function, the latter is merely a $C^{2}$-convex function [Ebe96, §1.6.6]. In the nonnegative sectional curvature situation for instance, a key idea in the fundamental work [CG72] is the construction of suitable continuous convex functions related to Busemann functions and having appropriate additional properties.
A function is (strictly) convex if its Hessian is positive (definite) semidefinite everywhere. If a Riemannian manifold admits a strictly convex function, then it has no compact immersed minimal submanifolds. In particular, it has no closed geodesics. When the strictly convex function has a minimum, such a manifold also allows for a generalisation of Cartan’s theorem [Yam82]: every compact subgroup of the isometry group has a common fixed point. The existence of strict convex functions also has consequences for the $\mathrm{L^{2}}$-spectrum of the Laplacian. The study of the spectrum of the Laplacian on complete Riemannian manifolds has a long, rich history and there are very many results, often with various additional constraints placed on the structure of the manifold and on its curvature. For instance, in the case of manifolds of nonpositive curvature the essential spectrum was shown to be an infinite interval depending on the dimension and sectional curvatures at infinity in [Don81, Theorem 6.3] using the distance function. One can then ask interesting questions about the finer structure of the spectrum e.g., whether there are eigenvalues in the essential spectrum. For a succinct recent overview of some of these aspects, we refer to [Don10]. We now content ourselves by mentioning a few results which use convexity in an essential manner to deduce refined information on the nature of the spectrum and to put our work in context, with no pretensions to completeness. In the case of nonnegative sectional curvature, explicit convex functions have also been used to show that there are no $\mathrm{L}^{2}$-eigenfunctions under the additional assumptions on the codimension of the soul of the manifold being $1$ or having a quadratic decay of radial curvatures when the soul is just a point [EF92, Theorems 4.2, 4.3]. On Hadamard manifolds, [Xav88, Theorem 1] exhibited a strictly convex function related to the square of the distance function satisfying the estimates required to ensure the nontriviality of the absolutely continuous spectrum. However, this is not sufficient to exclude the existence either of eigenvalues or of the singularly continuous spectrum.
In this article, we consider convex functions on complete Riemannian manifolds $M$ whose Jacobi fields satisfy certain growth properties, while making no assumptions on sectional curvatures. A simply connected Riemannian manifold is said to have no focal points if the norm of any nontrivial Jacobi field which vanishes at the initial point of a geodesic ray is strictly increasing. Since the exponential map at any point on such a manifold is a diffeomorphism, i.e. every point is a pole, they are necessarily noncompact. This class contains all Hadamard manifolds; however, it is well known [Gul75] that there are both compact and noncompact manifolds without focal points whose sectional curvatures change sign. While several aspects of the geometry of these manifolds have been well-studied, from understanding the dynamics of the geodesic flow [Ebe72], [O’S74], [Esc77]) to recent work on X-ray transforms on such manifolds with strictly convex boundary [PSUZ19] and the solution to the Dirichlet problem at infinity [LZ25], some aspects such as the spectrum of the Laplacian have not been analysed as much. We study some of these using the convexity of two kinds of geometrically defined functions on them. One of them is the square of the distance function from an arbitrary but fixed point. This is a smooth strictly convex function [Ebe72, Lemma 1] whose sub-level sets are compact. It is suitably modified to ensure a bounded gradient, an aspect that is pivotal in many applications. The other is the Busemann function associated to a unit vector in the tangent bundle of $M$. These are $C^{2}$ convex functions [Esc77, Theorem 2] but are not strictly convex. We remedy the situation by giving a new construction of a strictly convex function from Busemann functions. This leads to some new spectral theoretic results. A careful analysis of Jacobi fields and Jacobi tensors is central to these considerations. We now proceed to discuss our results in some detail.
Our first application using the strict convexity of the suitably modified distance function concerns the isoperimetric profile function $\mathcal{I}_{M}$ defined on the manifold $M$. This function maps every positive volume to the infimum of the perimeters of sets in $M$ having that volume. The isoperimetric profile is shown [Rit17, Theorem 3.1] to be continuous and nondecreasing for complete manifolds which possess smooth strictly convex Lipschitz exhaustion functions and is therefore differentiable almost everywhere on such manifolds. Examples are discussed under sectional curvature conditions there, namely, Hadamard manifolds and manifolds with strictly positive sectional curvatures. In contrast, we observe in Proposition 3.2 that for simply connected Riemannian manifolds without focal points the map $\mathcal{I}_{M}$ is a nondecreasing continuous function, thus giving a class of examples without constraints on sectional curvatures.
Strictly convex functions on Kähler manifolds can be seen to be strictly plurisubharmonic functions. So, manifolds without focal points which are also Kähler, admit strictly plurisubharmonic functions obtained, for instance, via the square of the distance function. From this, we observe in Proposition 3.4 that such manifolds are Stein using Grauert’s proof of Levi’s problem [GW77, §4]. By composing with a suitable strictly convex function on $\mathbb{R}$, we get a strictly convex exhaustion function with bounded gradient. This is applied in Proposition 3.5 to obtain a lower bound for the growth of the volume of metric balls following [NT03, Lemma 4.2 (i)]. The growth of the volume of the metric balls has been usually studied with lower bounds on Ricci curvature and upper bounds on sectional curvatures. It is an open problem whether a simply connected harmonic manifold which admits a Kähler structure is a symmetric space. Indeed in the Damek-Ricci setting this is known [BTV95, §4.1.12]. Therefore, at least in certain special cases such as harmonic manifolds without focal points, this volume growth is in line with the above mentioned rigidity by Bishop-Günther’s comparison theorem [Cha06, Theorem III.4.2].
We now proceed to describe our first spectral-theoretic result and note that the smooth strictly convex function with bounded gradient discussed in the previous paragraph is used here as well. A key idea is due to [Xav88] showing that on complete Riemannian manifolds supporting a sufficiently regular strictly convex function, Kato’s theory of $H$-smooth operators can be applied to understand the spectrum of the Laplacian. For a detailed discussion on $H$-smoothness we refer to the comprehensive survey [Sim19, §4]. In the case of Hadamard manifolds, this was done in [Xav88, Theorem 1] under suitable bounds on the Laplacian and its derivatives to conclude that the absolutely continuous part of the spectrum was nontrivial. In fact, it contains an interval that depends explicitly on these bounds. In the present work, we extend the implementation of $H$-smoothing to simply connected Riemannian manifolds without focal points in Theorem 3.8. Assuming only the nonpositivity of radial curvatures with respect to a point, a condition that is markedly weaker than global nonpositivity of sectional curvature, and bounds on Laplacian and the bilaplacian, we show that the absolutely continuous spectrum is an infinite interval which depends on the bounds. This is carried out by applying Lemma 3.6 where we obtain the required quantitative estimates on the Hessian of distance function in the usual fashion, while using a function that is defined when the manifold necessarily has no focal points. For context, we mention some of the work that derive information about the spectrum from the behaviour of the radial curvatures. Absence of point spectrum has been shown for manifolds with a pole satisfying quadratic decay of radial curvatures [Don10, Theorem 1.5]. Under linear decay of radial curvatures it was proved in [Kum10] that there are no eigenvalues embedded in the essential spectrum. If the radial curvature is assumed to approach a strictly negative constant near infinity, then [JL20] investigated the nature of the singular spectrum.
In view of Theorem 3.8, we can ask for more refined properties of the spectrum such as identifying whether the point spectrum or the singularly continuous spectrum are nontrivial. A sufficient condition for the absence of point spectrum and the singularly continuous spectrum is given by the existence of a strictly convex function whose bilaplacian is nonpositive [Xav88, Corollary to Theorem 2]. As noted there, this seemingly unnatural additional condition is, in fact, necessary. Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed. This sufficiency condition seems to us not to be approachable by the use of distance function.
To overcome this, we now turn our attention on the other natural geometrically defined convex function on simply connected Riemmanian manifolds without focal points, the Busemann function. Note that the Busemann function is convex but not strictly convex due to the vanishing of Hessian in the direction of its gradient. We use the geometry at infinity to construct a strictly convex function in Theorem 4.1 by averaging over all Busemann functions. This was motivated by the idea outlined in the context of symmetric spaces of noncompact type in [DX06, p-657], and is possibly of independent interest. For this, we make use of Jacobi tensors introduced in [Esc77] and its relation to the Hessian of Busemann functions for manifolds of continuous asymptote. In particular, for manifolds without focal points these Hessians can be seen to be positive semidefinite. The Hessian of the convex function which we construct by averaging over Buseman functions becomes positive definite under the assumption of negative Ricci curvature, yielding Theorem 4.1.
We now examine the hypotheses of Theorem 4.1. In this class of manifolds, having nonnegative Ricci curvature at even one point, forces the manifold to be flat [MZ00, Corollary 1]. Being an average of sectional curvatures, the sign of the Ricci curvature is a milder hypothesis. That this is true especially for negative Ricci curvature can be seen from the remarkable result [Loh94, Theorem A] that any smooth manifold of dimension at least three admits a complete Riemannian metric with negative Ricci curvature. Well-known classes of manifolds satisfying our assumptions include Riemannian symmetric spaces of noncompact type as well as simply connected noncompact (nonflat) harmonic manifolds without focal points. Harmonic manifolds are Einstein [Kni16] i.e, $Ric_{g}=\kappa g$, and have horospheres, which are level sets of Busemann functions, of constant mean curvature $h\geq 0$. If the Ricci curvature $\kappa>0$, then the harmonic manifold is necessarily a compact rank one symmetric space and if $\kappa=0$, then it is necessarily Euclidean. All other simply connected harmonic manifolds have negative Ricci curvature. Rank one symmetric spaces of noncompact type and other Damek-Ricci spaces are examples of this class of manifolds. We remark here that $h>0$ for such manifolds, since a harmonic manifold is Euclidean when $h=0$ [Kni16, §5, Theorem 19]. Harmonic manifolds without focal points are of purely exponential volume growth, and these have been well studied in the recent past (e.g.[Kni16, BKP21, Bra24]).
As principal applications of Theorem 4.1, we fully describe the nature of the spectrum $\sigma(\Delta)$ of the Laplacian $\Delta$ for interesting classes of manifolds having horospheres of constant mean curvature $h$. Simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$ are known as asymptotically harmonic manifolds [Kni16]. For all asymptotically harmonic manifolds, we show that equality holds in Cheeger’s inequality and that the Cheeger’s constant is equal to $h$ (see [KP15, §5] for related results). We mention in this context that for asymptotically harmonic Hadamard manifolds, the essential spectrum has been determined in [CS14, Theorem 4.4]. But it did not address finer questions such as existence of embedded eigenvalues or the singularly continuous spectrum. It was proved in [BMP23] that for asymptotically harmonic Hadamard manifolds with strictly negative sectional curvature (and so necessarily without focal points) the spectrum is absolutely continuous. On simply connected manifolds without focal points and negative Ricci curvature, as in Theorem 4.1, which are asymptotically harmonic, we show in Theorem 4.3 that the spectrum is purely absolutely continuous and is contained in $h^{2}/4,\infty)$.
Now since harmonic manifolds are asymptotically harmonic, the above result also holds in this case (Corollary 4.7). We give an elegant proof of the equality $\sigma(\Delta)=[h^{2}/4,\infty)$, making use of the Fourier transform developed recently in the context of harmonic manifolds without focal points in [BKP21]. It is not known whether this Fourier transform is unitary on the space of all $\mathrm{L}^{2}$-functions. However, the Fourier transform is unitary on radial $\mathrm{L}^{2}$-functions [BKP21, Theorem 4.7] with the Plancherel density given by the $\mathrm{c}$-function, and intertwines the radial Laplacian and the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$. Using certain estimates on the $\mathrm{c}$-function, we show that the essential range of the multiplication operator $M_{(\lambda^{2}+\frac{h^{2}}{4})}$ is equal to $h^{2}/4,\infty)$, enabling us to conclude the desired equality for the spectrum.
Although symmetric spaces of higher rank are not asymptotically harmonic, our geometric methods allow us to recover the absolute continuity of the spectrum. It is, of course, classical that the unitarity of the Fourier transform in the case of symmetric spaces of noncompact type completely determines the spectrum of the Laplacian to be absolutely continuous and equal to $\left\lVert\rho^{2}\right\rVert,\infty)$.
2. Preliminaries
In this section, we give a brief summary of the basic geometric setup needed in the sequel. Let $M$ be a complete Riemannian manifold and let $\gamma$ be a geodesic in $M$. A vector field $Y$ along $\gamma$ is called a Jacobi field if it satisfies the Jacobi equation
(1) $Y^{{}^{\prime\prime}}+R(\gamma^{{}^{\prime}},Y)\gamma^{{}^{\prime}}=0.$
Here $\gamma^{{}^{\prime}}$ is the tangent vector field of $\gamma$, $R$ is the Riemann curvature tensor on $M$ and $Y^{{}^{\prime}}$ denotes the covariant derivative of $Y$ with respect to $\gamma^{{}^{\prime}}$ along $\gamma$. The prototypical example of a Jacobi field is the variation vector field of a geodesic variation. Two points on $\gamma$ are said to be conjugate along $\gamma$ if there exists a nonzero Jacobi field which vanishes at both the points. $M$ is called a manifold without conjugate points if for every geodesic $\gamma$ in $M$, no two points on $\gamma$ are conjugate along $\gamma$ or equivalently, if the exponential map is non-singular at every point in $M$. In fact, $M$ has no conjugate points if and only if $\exp_{p}:T_{p}M\to M$ is a covering map for each $p$ in $M$. So if $M$ is also simply connected, then $\exp_{p}$ is a diffeomorphism for all $p\in M$ [O’S74, Theorem 1] and hence $M$ is necessarily noncompact. In this case, any two distinct points of $M$ are joined by a unique geodesic.
A simply connected Riemannian manifold $M$ is said to have no focal points if for every nontrivial Jacobi field $Y$ satisfying $Y(0)=0$, the map $t\mapsto\left\lVert Y(t)\right\rVert$ is strictly increasing for $t>0$ [O’S74, Proposition 4]. In particular, any non-trivial Jacobi field cannot vanish more than once and hence every manifold without focal points is also a manifold without conjugate points. Note that all Hadamard manifolds are manifolds without focal points.
For some applications, it is not just convenient but even important to consider the full vector space of normal Jacobi fields along a unit speed geodesic together rather than consider them one at a time. A Jacobi tensor is the tool which does this. We now give a brief overview of Jacobi tensors and refer to [Esc77] for a full account. Let $M$ be a simply connected Riemannian manifold without conjugate points and let ($SM$) TMTM denote its (unit) tangent bundle. For $v\in SM$, let $\gamma_{v}$ denote the unit speed geodesic determined by $v$. A normal $(1,1)$ - tensor $Y_{v}$ along the geodesic $\gamma_{v}$ is a bundle endomorphism of the normal bundle $N\gamma_{v}$ over $\gamma_{v}$ given by
$Y_{v}:{\mathbb{R}}\to\mathrm{End}N(\gamma_{v})=\bigcup\limits_{t\in{\mathbb{R}}}\mathrm{End}(N(\gamma_{v}(t))),\,N(\gamma_{v}(t)):=\left\{w\in T_{\gamma_{{}_{v}}(t)}M:{\langle w},{\gamma_{v}^{{}^{\prime}}(t)\rangle}_{\gamma_{{}_{v}}(t)}=0\right\}.$
A Jacobi tensor $Y_{v}$ is a normal $(1,1)$ - tensor field along a geodesic $\gamma_{v}$ with transversal derivative (i.e. there exists some $t$ such that $ker\penalty 10000\ Y_{v}(t)\cap ker\penalty 10000\ Y_{v}^{{}^{\prime}}(t)=0$) satisfying the Jacobi equation
(2) $Y_{v}^{{}^{\prime\prime}}(t)+R(t)Y_{v}(t)=0$
Here $R(t)$ is the symmetric $(1,1)$-tensor along $\gamma_{v}$ induced by the curvature tensor $R$ and defined by
$R(t)w:=R(\gamma_{v}^{{}^{\prime}}(t),w)\gamma_{v}^{{}^{\prime}}(t)$
and $Y_{v}^{{}^{\prime}}$ is the covariant derivative of the tensor $Y_{v}$ along $\gamma_{v}$ with respect to the tangent vector field $\gamma_{v}^{{}^{\prime}}$. For each $s>0$ and $v\in SM$, let $D_{s,v}$ be a solution to the Jacobi equation (2) with the boundary conditions $D_{s,v}(0)=\mathrm{Id}$ and $D_{s,v}(s)=0$. Then $D_{v}:=\lim\limits_{s\to\infty}D_{s,v}$ exists and is referred to as the stable Jacobi tensor $D_{v}$ satisfying $D_{v}(0)=\mathrm{Id}$. A simply connected Riemannian manifold without conjugate points is said to be a manifold of continuous asymptote if $v\to D_{v}^{{}^{\prime}}(0)$ is a continuous section of the vector bundle $\{(v,\phi):v\in SM,\phi\in\mathrm{End}(N(\gamma_{v})(0))\}$ over $SM$. As stated explicitly in [Got79, p-596] simply connected manifolds without focal points fall in this class. For such manifolds, [KP13, §11] gives a compactification known as the Busemann compactification. Let $o\in M$ be an arbitrary but fixed point. Then the Busemann boundary $\partial^{o}M$ is homeomorphic to $S_{p}M$ for any $p\in M$ [KP13, Proposition 11.4 (2)].
Convex functions being the theme of this article, the two natural geometrically defined functions which under suitable constraints on curvature are convex are the distance function and the Busemann function defined below. Given $t\in[0,\infty)$ and $v\in SM$, let $b_{t,v}:M\to\mathbb{R}$ denote the function defined by
$b_{t,v}(x):=d(x,\gamma_{v}(t))-t$
which, by triangle inequality, is increasing as a function of $t$. Since $b_{t,v}(\cdot)$ is absolutely bounded by $d(\gamma_{v}(0),\cdot)$, the function
$b_{v}:=\lim\limits_{t\to\infty}b_{t,v}$
is defined everywhere on $M$. The function $b_{v}$ is called the Busemann function associated to $v\in SM$. The level sets of Busemann functions are called horospheres. Horospheres are closed and hence are complete hypersurfaces with the induced metric. It is classical that [Ebe96, §1.6.6] for Hadamard manifolds the Busemann functions are $C^{2}$ convex functions. Eschenburg [Esc77] was able to relax the curvature restriction by making use of convergence properties of Jacobi tensors, and generalized the $C^{2}$ - regularity of horospheres to manifolds of continuous asymptote. In order to obtain this, the following relation between the Hessian of the Busemann function $b_{v}$ and the stable Jacobi tensor $D_{v}$ for each $v\in SM$ is used. We first make a small observation. The orthogonal projection $x^{\perp}(p,v)$ of $x\in T_{p}M$ onto $(\nabla b_{v}(p))^{\perp}$ will be denoted by $x^{\perp}$ for notational convenience.
Lemma 2.1.
For $p\in M$ and $x\in S_{p}M$, we have $\mathrm{Hess}\penalty 10000\ b_{v}(p)(x,x)=\mathrm{Hess}\penalty 10000\ b_{v}(p)(x^{\perp},x^{\perp})$.
Proof.
As $\nabla b_{v}$ is a $C^{1}$ vector field with $\left\lVert\nabla b_{v}\right\rVert=1$, we have that $\nabla_{\nabla b_{v}}\nabla b_{v}=0$. Indeed ${\langle\nabla_{X}\nabla b_{v}},{Y\rangle}=\mathrm{Hess}\penalty 10000\ b_{v}(X,Y)$ is symmetric in $X,Y$ and hence for every vector field $X$
${\langle\nabla_{\nabla b_{v}}\nabla b_{v}},{X\rangle}$ $={\langle\nabla_{X}\nabla b_{v}},{\nabla b_{v}\rangle}$
$=\frac{1}{2}X{\langle\nabla b_{v}},{\nabla b_{v}\rangle}=0$
In particular, we see that every integral curve of $\nabla b_{v}$ is a geodesic and hence complete. Then, the orthogonal decomposition $T_{p}M={\mathbb{R}}\nabla b_{v}\oplus(\nabla b_{v})^{\perp}$ enables us to write $x={\langle x},{\nabla b_{v}(p)\rangle}_{T_{p}M}\nabla b_{v}(p)+x^{\perp}$. Using $\nabla_{\nabla b_{v}}\nabla b_{v}=0$ and the bilinearity of $\mathrm{Hess}\penalty 10000\ b_{v}$ we have
$\mathrm{Hess}\penalty 10000\ b_{v}(p)(x,x)=\mathrm{Hess}\penalty 10000\ b_{v}(p)(x^{\perp},x^{\perp}).$
Notice in the above computation that when $X=\nabla b_{v}$ we have $\mathrm{Hess}\penalty 10000\ b_{v}(\nabla b_{v},\nabla b_{v})=0$ and hence $b_{v}$ is not strictly convex for any $v\in SM$. In fact, to the best of our knowledge, it is not known if Busemann functions are even convex for an arbitrary manifold of continuous asymptote. Using Lemma 2.1 and [Esc77, Proposition 2, Theorem 1] we have the following relation between the stable Jacobi tensors and the Busemann functions in the case of manifolds of continuous asymptote: for $v\in SM$, $p\in M$, and $x\in T_{p}M$,
(3) $\mathrm{Hess}\penalty 10000\ {b_{v}}(p)(x,x)=-{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)x^{\perp}},{D_{\nabla b_{v}(p)}(0)x^{\perp}\rangle}$
In what follows, we shall make use of the foregoing discussion for simply connected Riemannian manifolds without focal points and henceforth make this assumption.
3. A construction using the distance function and its applications
In this section, we consider the distance from a fixed point on a simply connected manifold $M$ without focal points, and the strictly convex functions derived from it, having additional properties required for our applications. For a fixed $p\in M$, the function $q\mapsto d^{2}(p,q)$ is smooth and strictly convex where $r(q)=d(p,q)$ is the distance function. This is proved by an analysis of Jacobi fields as follows. Note that the normalized distance-squared function $u(q)=\frac{1}{2}r^{2}(q)$ is smooth on $M$. The Hessian of $u$ is given by
$\mathrm{Hess}\penalty 10000\ u=\mathrm{d}r^{2}+r\mathrm{Hess}\penalty 10000\ r$
Given $v\in T_{p}M$, consider the Jacobi field along $t\mapsto\exp_{p}(tv)$ satisfying $Y(0)=0$. Then a routine calculation yields,
(4) $\mathrm{Hess}\penalty 10000\ r(\exp_{p}(tv))(Y(t),Y(t))={\langle Y^{{}^{\prime}}(t)},{Y(t)\rangle}=\frac{1}{2}\frac{d}{ds}\Big|_{s=t}\left\lVert Y(s)\right\rVert^{2}$
Since $M$ has no focal points, $d\exp_{p}$ is nonsingular at $tv$ for any $w\in T_{q}M$ there exists a Jacobi field $Y$ along $\gamma$ such that $Y(0)=0$ and $Y(t)=w$. Using [O’S74, Theorem 4] we see that $\mathrm{Hess}\penalty 10000\ r>0$ for $t>0$ and so $u$ is strictly convex on $M$ and we refer to [Ebe72] for more details. It then follows from [Yam82, Theorem A] that each compact subgroup of the isometry group of $M$ has a common fixed point. This generalizes the classical theorem of Cartan for Hadamard manifolds.
Now we construct, in the usual way, a smooth strictly convex function on $M$ whose gradient is bounded by taking $f(q)=\sqrt{1+2u(q)}$; then $\left\lVert\nabla f\right\rVert\leq 1$ and
$\mathrm{Hess}\penalty 10000\ f$ $=\frac{1}{(1+2u)^{\frac{3}{2}}}\mathrm{d}r^{2}+\sqrt{\frac{2u}{1+2u}}\mathrm{Hess}\penalty 10000\ r$
is positive definite as $\mathrm{Hess}\penalty 10000\ r$ is strictly positive showing that $f$ is strictly convex. A real-valued function $\varphi$ on a manifold is called an exhaustion function if all its sub-level sets $\{\varphi\leq c\}$ are compact. Notice that the sub-level sets of $f$ are given by
$f^{-1}(c)=\begin{cases}\emptyset,&\text{if}\penalty 10000\ c<1\\ \{p\},&\text{if}\penalty 10000\ c=1\\ B(p,c),&\text{if}\penalty 10000\ c>1\end{cases}$
and hence are all compact. Here $B(p,c)$ denotes the closed ball of radius $c$ centred at $p$. Thus $f$ is also an exhaustion function on $M$. For convenience, we summarize this discussion as a lemma.
Lemma 3.1.
Let $M$ be a simply connected Riemannian manifold without focal points. Then it possesses a smooth strictly convex exhaustion function with bounded gradient.
As a first application we have the following result on the isoperimetric profile function $\mathcal{I}_{M}$ on $M$. The isoperimetric profile of a Riemannian manifold is the function that assigns to a given positive volume the infimum of the perimeter of the sets of this volume. We remark here that regularity properties of this function are of interest and refer to the introduction in [Rit17] for a detailed discussion. The continuity and the nondecreasing nature of the isoperimetric profile is shown [Rit17, Theorem 3.1] for complete manifolds which possess smooth strictly convex Lipschitz exhaustion functions. It then follows from a theorem of Lebesgue that $\mathcal{I}_{M}$ is differentiable almost everywhere on such manifolds. As examples there, Hadamard manifolds and manifolds with strictly positive sectional curvatures are discussed. We give below a class of examples for which the continuity of $\mathcal{I}_{M}$ holds while making no assumptions on curvature and which properly contains Hadamard manifolds. As noted before this class contains plenty of examples whose sectional curvatures do not have a fixed sign [Gul75].
Proposition 3.2.
Simply connected Riemannian manifolds without focal points have a continuous nondecreasing isoperimetric profile function. In particular, they are differentiable almost everywhere.
Proof.
By Lemma 3.1, $M$ admits a smooth strictly convex Lipschitz exhaustion function. Invoking [Rit17, Theorem 3.1], we see that $\mathcal{I}_{M}$ is nondecreasing and continuous. ∎
Regularity for $\mathcal{I}_{M}$ has many interesting consequences; for instance, we have the following result whose proof makes use of standard arguments.
Proposition 3.3.
Let $M$ be as in Proposition 3.2. Suppose $v_{0}$ is a differentiable point of $\mathcal{I}_{M}$. If $\Omega$ is an isoperimetric region in $M$ of volume $v_{0}$, then (the regular part of) its boundary $\partial\Omega$ has constant mean curvature $H=\mathcal{I}_{M}^{{}^{\prime}}(v_{0})$.
Proof.
Since $\Omega$ is isoperimetric region the (regular part of) $\partial\Omega$ has constant mean curvature $H$. Let $\mu,\nu$ be the outward and inward normal variational vector fields on $\partial\Omega$ with compact support. Then the first variations of the volumes and the boundary areas are as follows
$\frac{d}{dt}V(\Omega_{t})\Big|_{t=0}$ $=\int\limits_{\partial\Omega_{0}}|\mu|\mathrm{d}A,\quad\frac{d}{dt}V(\Omega_{t})\Big|_{t=0}=-\int\limits_{\partial\Omega_{0}}|\nu|\mathrm{d}A$
$\frac{d}{dt}A(\partial\Omega_{t})\Big|_{t=0}$ $=H\int\limits_{\partial\Omega_{0}}|\mu|\mathrm{d}A,\quad\frac{d}{dt}A(\partial\Omega_{t})\Big|_{t=0}=-H\int\limits_{\partial\Omega_{0}}|\nu|\mathrm{d}A$
Hence, using the differentiability assumption at $v_{0}$,
$\lim\limits_{\delta v\to 0^{+}}\frac{\mathcal{I}_{M}(v_{0}+\delta v)-\mathcal{I}_{M}(v_{0})}{\delta v}\leq H,\quad\lim\limits_{\delta v\to 0^{-}}\frac{\mathcal{I}_{M}(v_{0}+\delta v)-\mathcal{I}_{M}(v_{0})}{\delta v}\geq H$
yields $\mathcal{I}_{M}^{{}^{\prime}}(v_{0})=H$. ∎
As a consequence, all isoperimetric regions in $M$ of the same volume $v_{0}$ must also have the same mean curvature if $v_{0}$ is a differentiable point of the isoperimetric profile function. In general, the existence of isoperimetric domains of a given volume in a complete noncompact Riemannian manifold is a difficult problem (see [Rit23]).
We now discuss a couple of direct applications in the situation when $M$ is, in addition, Kähler. For general facts on Kähler manifolds we refer to [GW77, §7]. On a Kähler manifold the complex analogue of convex functions is given by the notion of plurisubharmonic functions. These are defined using the complex Hessian called the Levi form. In local coordinates, the Levi form of a function $\tilde{f}$ is given by
$L\tilde{f}=4\sum\limits_{i,j=1}^{n}\frac{\partial^{2}h}{\partial z_{i}\partial\overline{z}_{j}}\mathrm{d}z_{i}\otimes\mathrm{d}\overline{z}_{j}$
where $(z_{1},\cdots,z_{n})$ is a local holomorphic coordinate system. Let $(x_{1},y_{1},\cdots,x_{n},y_{n})$ be the corresponding normal coordinates with $z_{k}=x_{k}+iy_{k}$, $k=1,\cdots,n$. A real-valued function on a Kähler manifold is said to be (strictly) plurisubharmonic, if the Levi form is positive (definite) semi-definite. Since $M$ is Kähler, the Levi form $L_{\tilde{f}}$ of $\tilde{f}$ is expressed in terms of Hessian of $\tilde{f}$ using the real coordinates as
$L_{\tilde{f}}=\mathrm{Hess}(\tilde{f})\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{1}}\right)+\mathrm{Hess}(\tilde{f})\left(\frac{\partial}{\partial y_{1}},\frac{\partial}{\partial y_{1}}\right).$
Therefore if $\tilde{f}$ is a strictly convex $C^{2}$-function on a Kähler manifold, then it is strictly plurisubharmonic.
Stein manifolds are complex manifolds which are biholomorphic to a closed complex submanifold of $\mathbb{C}^{n}$ for some $n\in\mathbb{N}$. This tells us that there are plenty of holomorphic functions on such manifolds. Stein manifolds are necessarily Kähler and a question of great interest then is to ask which Kähler manifolds are Stein. Typically, sufficient conditions are given by putting constraints on the sign of of sectional, holomorphic bisectional or Ricci curvatures[GW77, Theorem $1$]. It is in this context that we observe the following sufficient condition where we make no assumptions on curvature.
Proposition 3.4.
Let $M$ be a simply connected Riemannian manifold without focal points which is also Kähler. Then $M$ is necessarily Stein.
Proof.
Lemma 3.1 gives a smooth strictly convex exhaustion function on $M$, while the discussion above on the Levi form shows that this function is a strictly plurisubharmonic exhaustion function on $M$. Then Grauert’s proof of Levi’s problem [GW77, §4] that a complex manifold is Stein if and only if it admits a strictly plurisubharmonic exhaustion function concludes the proof. ∎
As another application of Lemma 3.1, we have the following volume growth comparison for metric balls in Kähler manifolds without focal points using [NT03, Lemma 4.2 (i)].
Proposition 3.5.
Let $M$ be a simply connected Riemannian manifold without focal points which is also Kähler and of (complex) dimension $m$. Then $M$ admits a smooth strictly plurisubharmonic function with bounded gradient. Moreover, for each $p\in M$ there exists a constant $C_{2}>0$ such that,
$V_{p}(r)\geq C_{2}r^{m},\quad r\geq 1$
where $V_{p}(r)$ is the volume of a metric ball of radius $r$ about $p$ in $M$.
As mentioned in the introduction, the question of whether a simply connected harmonic manifold that carries a Kähler structure must be a symmetric space is still unresolved. However, in the context of Damek–Ricci spaces this has been verified. Therefore, at least in certain special cases such as harmonic manifolds without focal points, the resulting volume growth reflects the same rigidity, in accordance with Bishop–Günther’s comparison theorem [Cha06, Theorem III.4.2].
Our next application is spectral theoretic in nature, and so we begin with a sketch of the basic functional analytic setup concerning the spectrum of the Laplacian referring to [Bor20] for the details. Let $M$ be a complete Riemannian manifold. Then the (positive) Laplacian $\Delta$ is an essentially self-adjoint operator with the space of compactly supported smooth functions on $M$ as the domain. We also denote its unique self-adjoint extension, having the domain $\mathcal{D}(\Delta):=\{f\in L^{2}(M):\Delta f\in L^{2}(M)\}$, by $\Delta$. Using the functional calculus for $\Delta$ we have the family of spectral projection operators $\prod_{E}:=\chi_{E}(\Delta)$ for each Borel subset $E$ of $\mathbb{R}$} referred to as the resolution of $\Delta$. The support of these projections and the spectrum of $\Delta$ are related in the following way. We say that $\lambda\in{\mathbb{C}}$ is in the spectrum $\sigma(\Delta)$ of $\Delta$ if and only if $\prod_{(\lambda-\epsilon,\lambda+\epsilon)}\neq 0$ for all $\epsilon>0$. If $\prod_{\{\lambda\}}\neq 0$, then $\lambda$ is an eigenvalue of $\Delta$ and the range of $\prod_{\{\lambda\}}$ is the corresponding eigenspace. The spectral theorem gives a unitary equivalence between $\Delta$ and a multiplication operator on a suitable Hilbert space of the form $\mathrm{L}^{2}(X,\mu)$, where $\mu$ is a $\sigma$- finite Borel measure on a measure space $X$. By the Lebesgue decomposition, $\mu$ decomposes as $\mu=\mu_{pp}+\mu_{ac}+\mu_{sc}$ into pure point measure, absolutely continuous and singularly continuous measures with respect to the Lebesgue measure, and consequently we have that
(5) $\mathrm{L}^{2}(X,\mu)=\mathrm{L}^{2}(X,\mu_{pp})\oplus\mathrm{L}^{2}(X,\mu_{ac})\oplus\mathrm{L}^{2}(X,\mu_{sc}).$
This induces a decomposition $\mathrm{L}^{2}(M)=\mathcal{H}_{pp}\oplus\mathcal{H}_{ac}\oplus\mathcal{H}_{sc}$ from which we have the associated decomposition of the spectrum of $\Delta$ namely, $\sigma(\Delta)=\overline{\sigma_{pt}(\Delta)}\cup\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ where
$\sigma_{ac}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{ac}}}),\quad\sigma_{sc}(\Delta)=\sigma(\Delta_{|_{\mathcal{H}_{sc}}}),\quad\overline{\sigma_{pt}(\Delta)}=\sigma(\Delta_{|_{\mathcal{H}_{pp}}})$
Here we take the closure for the set of all eigenvalues as it is not necessarily closed. We refer to $\sigma_{ac}(\Delta)$ and $\sigma_{sc}(\Delta)$ as the absolutely continuous and singularly continuous parts of $\sigma(\Delta)$ respectively and their union as the continuous part $\sigma_{cont}(\Delta)$ of $\sigma(\Delta)$. There is also a decomposition of $\sigma(\Delta)$ as a disjoint union:
$\sigma(\Delta)=\sigma_{disc}(\Delta)\cup\sigma_{ess}(\Delta)$
where $\sigma_{disc}(\Delta)$ and $\sigma_{ess}(\Delta)$ denote the discrete and essential parts of $\sigma(\Delta)$. The former consists of all eigenvalues of finite multiplicities and hence is contained in $\sigma_{pt}(\Delta)$. The latter is the complement of $\sigma_{disc}(\Delta)$ in $\sigma(\Delta)$ thereby we have that the continuous part of the spectrum, namely $\sigma_{ac}(\Delta)\cup\sigma_{sc}(\Delta)$ is contained in $\sigma_{ess}(\Delta)$. Note that it is possible that $\sigma_{ess}(\Delta)$ contains eigenvalues of infinite multiplicities usually referred to as the embedded eigenvalues.
In order to analyse the nature of the spectrum of the Laplacian using convexity, we make use of [Xav88, Theorem 3]. The idea there was to use Kato’s $H$-smooth operators to show that if $M$ admits a strictly convex function, then certain boundedness conditions on its gradient and Laplacian implies the non-triviality of the absolutely continuous part of the spectrum. This is then applied in the context of Hadamard manifolds with certain additional assumptions which are satisfied with lower bounds on sectional curvatures, to show the nontriviality of the absolutely continuous spectrum. It is this result that we now generalise to simply connected Riemannian manifolds without focal points having nonpositive radial curvature with respect to a single point. This is a much weaker condition than the nonpositivity of sectional curvature. To this end, we prove the following lemma.
Lemma 3.6.
Let $M$ be a simply connected Riemannian manifold without focal points. Suppose there exists a point $p$ in $M$ such that the radial curvatures with respect to $p$ are nonpositive. Then we have for any $q\neq p$ in $M$
$r(q)\mathrm{Hess}\penalty 10000\ r(q)(v,v)\geq\left\lVert v\right\rVert^{2}\qquad\forall\penalty 10000\ v\in T_{q}M$
Proof.
Since $M$ has no focal points $\exp_{p}:T_{p}M\to M$ is a diffeomorphism. Hence for any $q\in M$, $q\neq p$, there exists a unique $t\in{\mathbb{R}}$ and a unit tangent vector $v\in T_{p}M$ such that $q=\exp_{p}(tv)$. Moreover, $\gamma(s)=\exp_{p}(sv)$ is a unit speed geodesic in $M$ joining $p$ and $q$ satisfying $\gamma(0)=p$ and $\gamma(t)=q$. Then $\gamma^{\prime}(s)=\nabla r(\gamma(s))$ and in particular, $\gamma^{\prime}(t)=\nabla r(q)$. Since $d\exp_{p}$ is nonsingular at $tv$, for any $w\in T_{q}M$ there exists a Jacobi field $Y$ along $\gamma$ such that $Y(0)=0$ and $Y(t)=w$. Notice that by our assumption on nonpositive radial curvature, we have ${\langle R(Y(t),\gamma^{{}^{\prime}}(t))\gamma^{{}^{\prime}}(t)},{Y(t)\rangle}\leq 0$ for any $t>0$ and $v\in S_{p}M$. We now recall the standard technique of Jacobi field comparison. Since the second derivative of the Jacobi field is related to the curvature tensors via the Jacobi equation, quantitative estimates are obtained by imposing conditions on curvatures. Note that
$\lambda(s)=\frac{\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}},\penalty 10000\ s>0$
is well defined since $M$ has no focal points. Using l’Hospital rule, we see that $\lambda(0)=0$ since $Y^{{}^{\prime}}(0)\neq 0$. Differentiating and using the Jacobi equation we have
$\lambda^{{}^{\prime}}(s)$ $=\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-\left(\left\lVert Y^{{}^{\prime}}(s)\right\rVert^{2}-{\langle R(\gamma^{{}^{\prime}}(s),Y(s))\gamma^{{}^{\prime}}(s)},{Y(s)\rangle}\right)\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}$
$\leq\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-\left\lVert Y^{{}^{\prime}}(s)\right\rVert^{2}\left\lVert Y(s)\right\rVert^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}\quad\text{(nonpositivity of radial curvature)}$
$\leq\frac{2{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}-{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}{{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}^{2}}=1\quad\text{(Cauchy-Schwarz inequality)}$
Hence $\lambda(s)\leq s$ or equivalently,
$s{\langle Y^{{}^{\prime}}(s)},{Y(s)\rangle}\geq\left\lVert Y(s)\right\rVert^{2},\quad s>0$
The claim follows from (4). ∎
We are now ready to prove the main spectral result of this section. The bounds on the Laplacian and the bilaplacian demanded below are satisfied if lower bounds on curvatures and derivatives of metric tensor are assumed. It should be kept in mind that there exist Hadamard manifolds with pure point spectrum if the curvatures become unbounded.
Remark 3.7.
Note that we adopt the convention of considering the Laplacian as a positive operator, in contrast to the convention used in [Xav88] for Hadamard manifolds. Consequently, certain inequalities in the subsequent theorem appear with opposite signs when compared to the corresponding proof in [Xav88].
Theorem 3.8.
Let $M$ be a simply connected Riemannian manifold without focal points. Suppose there exists a point $p$ in $M$ such that the radial curvatures with respect to $p$ are nonpositive. If there exist constants $c_{1},c_{2}>0$ such that $f$ and the Laplacian $\Delta$ satisfies
  1. (i)
    $\Delta f\geq-c_{1}$
  2. (ii)
    $\Delta^{2}f\leq c_{2}f^{-3}$
then the absolutely continuous spectrum of $\Delta$ is nontrivial. In fact, $(\alpha,\infty)$ is in the absolutely continuous part of the spectrum of $\Delta$ where $\alpha=\frac{3}{2}c_{1}+\frac{1}{4}c_{2}$
Proof.
Note that $\left\lVert\nabla f\right\rVert$ is bounded. Since $M$ has nonpositive radial curvature at $p$, from Lemma 3.6 we obtain that $\mathrm{Hess}\penalty 10000\ r^{2}(p)$ is the Riemannian metric at $p$ and for any $p\neq q\in M$,
$\mathrm{Hess}\penalty 10000\ r^{2}(q)(w,w)={\langle\nabla r(q)},{w\rangle}^{2}+r(q)\mathrm{Hess}\penalty 10000\ r(q)(w,w)\geq\left\lVert w\right\rVert^{2},\quad\forall\penalty 10000\ w\in T_{q}M,\left\lVert w\right\rVert=1$
From here, the arguments are standard, as in [Xav88, Theorem 1], for example. We write them down here for completeness. Recall that $f(q)=(1+r(q)^{2})^{1/2}$ and its Hessian can be computed to be
$\mathrm{Hess}\penalty 10000\ f(q)(w,w)$ $=-\frac{r^{2}(q)}{(1+r^{2}(q))^{3/2}}{\langle w},{\nabla r(q)\rangle}^{2}+\frac{1}{2(1+r^{2}(q))^{1/2}}\mathrm{Hess}\penalty 10000\ r^{2}(q)(w,w)$
$\geq\frac{1}{(1+r^{2}(q))^{3/2}}$
Taking $h(q)=\frac{1}{(1+r^{2}(q))^{3/2}}$, we see that
$\Delta h$ $=\frac{3}{f^{4}}\Delta f+\frac{12}{f^{5}}\left\lVert\nabla f\right\rVert^{2},$
$\implies\Delta h$ $\geq\frac{3\Delta f}{f^{4}}\geq-\frac{3c_{1}}{2}$
Thus for
$\beta=\inf\frac{1}{4h}\{2\Delta h-\Delta^{2}f\}\geq-\frac{3c_{1}}{2}-\frac{c_{2}}{4}$
and the claim now follows as in [Xav88, Theorem 3]. ∎
Clearly, the above theorem generalizes [Xav88, Theorem 1] under much weaker curvature conditions. However, more refined questions such as identifying whether the point spectrum or the singularly continuous spectrum occur are not amenable to this approach via the use of the distance function. To overcome this, we now turn our attention on the other natural geometrically defined convex function, the Busemann function.
4. A construction using Busemann functions and its applications
In this section, we construct a smooth strictly convex function out of Busemann functions on simply connected manifolds without focal points, using the geometry at infinity and assuming, in addition, strict negativity of Ricci curvature. As remarked in the introduction, this is motivated by a similar construction on symmetric spaces of noncompact type outlined in [DX06, penultimate paragraph p.657657]. The main purpose of this construction here is that it allows us to prove new results (Theorem 4.3 and Corollary 4.7) completely determining the nature of the spectrum for certain interesting classes of manifolds without focal points whose horospheres have constant mean curvature.
[…]
Theorem 4.1.
Let $M$ be simply connected Riemannian manifold with no focal points and negative Ricci curvature. Then $F(p)=\displaystyle{\int\limits_{M(\infty)}b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)}$ is a strictly convex function whose gradient is bounded.
Proof.
Let $o$ be an arbitrary but fixed point in $M$. Let $\{v_{k}\}$ be a countable dense subset in $S_{o}M$. Consider the Dirac measures $\delta_{v_{k}}$ associated to each $v_{k}$. Then ${\eta=\sum\limits_{k=1}^{\infty}\frac{1}{2^{k}}\delta_{v_{k}}}$ defines a probability measure on $S_{o}M$. We identify $S_{o}M$, $M(\infty)$ and $\partial^{o}M$ via the respective homeomorphisms. By dominated convergence theorem, the gradient of $F$ at $p$ is given by $\nabla F(p)=\displaystyle{\int\limits_{M(\infty)}\nabla b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)}$. Clearly,
$\left\lVert\nabla F(p)\right\rVert=\left\lVert\int\limits_{M(\infty)}\nabla b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)\right\rVert\leq\int\limits_{M(\infty)}\left\lVert\nabla b_{v}(p)\right\rVert\penalty 10000\ \mathrm{d}\eta(v)=1,\quad\text{as}\penalty 10000\ \left\lVert\nabla b_{v}\right\rVert=1$
i.e., $F$ has bounded gradient. In particular, $F$ is Lipschitz with Lipschitz constant $1$. For $p\in M$ and $w\in S_{p}M$, using (3) and that $\mathrm{Hess}\penalty 10000\ b_{v}$ is positive semidefinite we have
$\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v)$ $=\left|\int\limits_{M(\infty)}{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}\penalty 10000\ \mathrm{d}\eta(v)\right|$
$\leq\int\limits_{M(\infty)}|{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}|\penalty 10000\ \mathrm{d}\eta(v)$
$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w\right\rVert\left\lVert w\right\rVert\penalty 10000\ \mathrm{d}\eta(v)\quad\text{(using Cauchy-Schwarz)}$
$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\penalty 10000\ \mathrm{d}\eta(v)\quad(D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\penalty 10000\ \text{is a bounded linear map for}\penalty 10000\ v\in S_{p}M).$
The restriction of the continuous map (§2) $w\mapsto\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert$ on $SM$ to $S_{p}M$ shows that there exists some $L>0$ such that $\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $w\in S_{p}M$. In particular, $\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $v\in S_{o}M$. So another application of the dominated convergence theorem shows that
$\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)=\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v).$
The convexity of $F$ follows from the convexity of the Busemann functions. We now claim that $F$ is, in fact, strictly convex. Suppose there exists $p\in M$ and $0\neq w\in T_{p}M$ such that $\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)=0$ then
$\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v)=0$
Since $v\mapsto\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)$ is a nonnegative continuous function in $v$, we have $\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)=0$ for all $v\in S_{o}M$. Using (3),
(6) ${\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}},{D_{\nabla b_{v}(p)}(0)w^{\perp}\rangle}=0,\quad\text{for all}\penalty 10000\ v\in S_{o}M.$
It is known from [Esc77, §5] that for $v\in SM$ and all parallel vector fields $x$ normal to $\gamma_{v}$, $\frac{d}{dt}(\left\lVert D_{v}(t)x\right\rVert^{2})\leq 0$ i.e., the function $f_{v,x}(t)=\frac{1}{2}\left\lVert D_{v}(t)x\right\rVert^{2}$ is $C^{2}$ and monotonically decreasing. From (6) we see that for each $v\in S_{o}M$, $f_{\nabla b_{v}(p),w^{\perp}}$ has a critical point at 0. We claim that $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. If not, then 0 is either a local minimum or local maximum. Suppose 0 is a local minimum then there exists a $\delta$-neighbourhood of 0 such that $f_{\nabla b_{v}(p),w^{\perp}}(0)\leq f_{\nabla b_{v}(p),w^{\perp}}(t)$ for all $t\in(-\delta,\delta)$. Since $f$ is monotonically decreasing we have $f_{\nabla b_{v}(p),w^{\perp}}(t)=f_{\nabla b_{v}(p),w^{\perp}}(0)$ for all $t\in(0,\delta)$. Hence $f_{\nabla b_{v}(p),w^{\perp}}$ is constant on $(0,\delta)$. Observe that $f_{\nabla b_{v}(p),w^{\perp}}$ is $C^{2}$ and so $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=\lim\limits_{t\to 0_{+}}f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(t)=0$, a contradiction. A similar argument could be carried out to show that 0 cannot be a local maximum of $f_{\nabla b_{v}(p),w^{\perp}}$. Hence 0 is a degenerate critical point of $f_{\nabla b_{v}(p),w^{\perp}}$ i.e, $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. Then the Jacobi equation for $D_{\nabla b_{v}(p)}$,
$D_{\nabla b_{v}(p)}^{{}^{\prime\prime}}(0)+R_{\nabla b_{v}(p)}(0)D_{\nabla b_{v}(p)}(0)=0$
gives
$0$ $=f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)$
$={\langle D_{\nabla b_{v}(p)}^{{}^{\prime\prime}}(0)w^{\perp}},{w^{\perp}\rangle}+\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
$=-{\langle{R_{\nabla b_{v}(p)}}(0)w^{\perp}},{w^{\perp}\rangle}+\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
$\implies{\langle R_{\nabla b_{v}(p)}(0)w^{\perp}},{w^{\perp}\rangle}$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}$
(7) $\implies K(\nabla b_{v}(p),w^{\perp})$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}\geq 0,\quad\text{for all $v\in S_{o}M$}$
where $K$ denotes the sectional curvature corresponding to the plane spanned by $\nabla b_{v}(p)$ and $w^{\perp}$. Let $x\in S_{p}M$ be such that $x\perp w$. Since $S_{p}M$ is homeomorphic to $M(\infty)$ there exists a unique $v_{o}\in S_{o}M$ such that $\gamma_{v_{o}}$ and $\gamma_{x}$ are asymptotic. As $\{v_{k}\}$ is dense in $S_{o}M$ there exists a subsequence $(v_{k_{m}})$ in $S_{o}M$ converging to $v_{o}$. For each $m$ we have $\nabla b_{v_{{}_{k_{m}}}}(p)\in S_{p}M$ such that $\gamma_{\nabla b_{v_{{}_{k_{m}}}}(p)}$ and $\gamma_{v_{{}_{k_{m}}}}$ are asymptotic and so $v_{{}_{k_{m}}}$ and $\nabla b_{v_{{}_{k_{m}}}}(p)$ correspond to the same element in $M(\infty)$. It then follows that the sequence $(\nabla b_{v_{{}_{k_{m}}}}(p))$ converges to $x$ in $S_{p}M$. We denote the orthogonal projection of $w$ onto $(\nabla b_{v_{{}_{k_{m}}}}(p))^{\perp}$ by $w^{\perp}(m)$. Then
$\left\lVert w-w^{\perp}(m)\right\rVert^{2}$ $=\left\lVert{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}\nabla b_{v_{k_{m}}}(p)\right\rVert^{2}$
$=|{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}|^{2}\left\lVert\nabla b_{v_{k_{m}}}(p)\right\rVert^{2}$
$=|{\langle w},{\nabla b_{v_{k_{m}}}(p)\rangle}|^{2}$
$\xrightarrow{m\to\infty}0\quad\text{as $\nabla b_{v_{k_{m}}}(p)\to x$ and $x\perp w$}.$
Thus the sequence $(w^{\perp}(m))$ converges to $w$ as $m\to\infty$ and from (4) we have,
[…]
Now for $\{x_{1},\cdots,x_{n-1}\}$ orthonormal vectors in $T_{p}M$ such that $x_{i}\perp w$, $Ricc(w,w)=\sum\limits_{i=1}^{n}K(x_{i},w)\geq 0$. This is a contradiction to the hypothesis that $M$ has negative Ricci curvature. Hence $\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)>0$ for every $p\in M$ and every $0\neq w\in T_{p}M$ proving the strict convexity of $F$. ∎
As indicated at the beginning of the section, we will be concerned with the spectrum on asymptotically harmonic manifolds. Recall that these are simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$. In particular, we answer more refined questions of whether there is a nontrivial point spectrum or singularly continuous spectrum on these manifolds. The main tool here is an implementation of Kato’s theory of $H$-smooth operators on complete Riemannian manifolds due to [Xav88, Theorem 2]. When it can be carried out, it enables one to exclude the existence of the point spectrum and the singularly continuous spectrum. As noted in [Xav88], the seemingly unnatural additional condition $\Delta^{2}F\leq 0$ that is demanded is, in fact, necessary. Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed there. We shall observe below that the condition on the bilaplacian is satisfied for $M$ as in Theorem 4.1 which are in addition asymptotically harmonic.
Lemma 4.2.
Let $M$ and $F$ be as in Theorem 4.1. Further, assume that $M$ is asymptotically harmonic with $h$ as the constant mean curvature of the horospheres. Then $F$ has bounded Laplacian and $\Delta^{2}F=0$.
Proof.
Since $\Delta$ is the trace of Hessian, by choosing an orthonormal basis $\{E_{i}(p)\}_{i=1}^{n}$ of $T_{p}M$ we have
$\sum\limits_{i=1}^{n}\mathrm{Hess}\penalty 10000\ ^{F}(p)(E_{i}(p),E_{i}(p))$ $=\sum\limits_{i=1}^{n}\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(E_{i}(p),E_{i}(p))\penalty 10000\ \mathrm{d}\eta(v)$
$=\int\limits_{M(\infty)}\sum\limits_{i=1}^{n}\mathrm{Hess}\penalty 10000\ b_{v}(p)(E_{i}(p),E_{i}(p))\penalty 10000\ \mathrm{d}\eta(v)$
$\Delta F(p)$ $=\int\limits_{M(\infty)}\Delta b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)=h$
and the lemma follows. ∎
As an immediate consequence we have the following theorem applying [Xav88, Theorem 2].
Theorem 4.3.
If $M$ as in Theorem 4.1 is also asymptotically harmonic with $h$ as the constant mean curvature of the horospheres, then the spectrum of the Laplacian is purely absolutely continuous. Moreover, the spectrum is contained in $\frac{h^{2}}{4},\infty)$.
Proof.
The absolute continuity of the spectrum follows by combining [Xav88, Theorem 2] with Lemma 4.2. That the spectrum of Laplacian is contained in $\frac{h^{2}}{4},\infty)$ follows from the discussion below. ∎
We first briefly discuss some facts regarding the bottom $\lambda_{0}(M)$ of the spectrum for general asymptotically harmonic manifolds. For a complete noncompact Riemannian manifold $M$ it is characterized by [Cha84, §4.4]
(8) $\lambda_{0}(M)=\inf\limits_{f\neq 0}\frac{\int\limits_{M}|\nabla f|^{2}}{\int\limits_{M}|f|^{2}}$
where $f$ varies over the completion $\mathfrak{H}(M)$ of $C_{c}^{\infty}(M)$ with respect to the norm $\left\lVert f\right\rVert=\left\lVert f\right\rVert_{\rm{L}^{2}(M)}+\left\lVert\nabla f\right\rVert_{\rm{L}^{2}(M)}$. Let $|\Omega|$ denote the volume of a submanifold $\Omega$ in $M$. Then the Cheeger’s constant of $M$ is given by
$h_{Cheeger}(M):=\inf\limits_{\Omega\subset M}\frac{|\partial\Omega|}{|\Omega|}$
where the infimum varies over all open, connected submanifolds $\Omega$ of $M$ with compact closure and smooth boundary $\partial\Omega$. The importance of this global geometric invariant is illustrated by the following fundamental inequality relating it to the global analytic invariant $\lambda_{0}(M)$ [Cha84, §4.3, Theorem 3],
(9) $\frac{h_{Cheeger}^{2}(M)}{4}\leq\lambda_{0}(M).$
Proposition 4.4.
Let $M$ be an asymptotically harmonic manifold with horospheres of mean curvature $h$ i.e., $\Delta b_{v}=h$ for all $v\in SM$. Then we have $h_{Cheeger}=h$ and $\lambda_{0}(M)=h^{2}/4$.
Proof.
For an open connected $\Omega$ in $M$ with compact closure and smooth boundary $\partial\Omega$ we have, by Stokes’ theorem, that
$h|\Omega|=\int\limits_{\Omega}\Delta b_{v}=\left|\int\limits_{\partial\Omega}{\langle\nabla b_{v}},{\nu\rangle}\right|\leq|\partial\Omega|,\quad(\text{since $\left\lVert\nabla b_{v}\right\rVert=1$})$
where $\nu$ is the outer unit normal vector field of $\Omega$ along $\partial\Omega$. Thus,
(10) $h\leq\inf\limits_{\Omega\subseteq M}\frac{|\partial\Omega|}{|\Omega|}=h_{Cheeger}$
where the infimum varies over $\Omega$ as above in $M$. Let $o\in M$ be arbitrary. Since $M$ is simply connected without conjugate points the exponential at $o$ is a diffeomorphism. Let $d(o,\cdot):M\to{\mathbb{R}}$ denote the distance from $o$. By taking the sequence of functions $f_{n}(x)=e^{-\frac{1}{2}(h+\frac{1}{n})d(o,x)}$ in $\mathfrak{H}(M)$, $n\in{\mathbb{N}}$, we see from (8) that
(11) $\lambda_{0}(M)\leq\frac{h^{2}}{4}$
Combining (9), (10), (11) we have
(12) $h=h_{Cheeger}\quad\text{and}\quad\lambda_{0}(M)=\frac{h^{2}}{4}$
Remark 4.5.
We remark here that for the proposition above to hold it suffices that, in the case of simply connected Riemannian manifold without conjugate points, there is a single vector $v\in SM$ such that the corresponding horospheres have constant mean curvature.
We have the following immediate corollary to Theorem 4.3 following [Don81] when the sectional curvature is strictly negative.
Corollary 4.6.
Let $M$ be an asymptotically harmonic Hadamard manifold with strictly negative sectional curvature. Then the spectrum of the Laplacian is purely absolutely continuous and is equal to $\frac{h^{2}}{4},\infty)$.
Observe that for such manifolds the essential spectrum and the absolutely continuous spectrum have been determined in [CS14, Theorem 4.4] and [BMP23, Theorem 1.3] respectively.
Harmonic manifolds are important examples of asymptotically harmonic manifolds. We describe only the aspects of harmonic manifolds which are needed in the sequel and to set notation, and we refer to [Bes78, Chapter $6$], and to the comprehensive exposition [Kni16] for the details. A function $f$ on a Riemannian manifold $(M,g)$ is said to be radial around a point $o$ if $f$ is constant on geodesic spheres $S_{o}(r)$ of radius $r$ centered at $o$. $M$ is said to be harmonic if the volume density in normal coordinates centered at $o$ is a radial function. Given a continuous function $f$, its radialisation about $o$ given by
$R_{o}(f)(z):=\int\limits_{S_{o}(r)}f(y){\rm{d}}\sigma^{r}(y)$
where $\sigma^{r}$ denotes surface area measure on $S_{o}(r)$ (induced from the Riemannian metric on $M$), normalized to have mass one. Note that for $f$ radial around $o$, we have that $R_{o}f=f$. Harmonicity of the manifold can be seen to be equivalent to the fact that the radialisation operator $R_{o}$ commutes with the Laplacian $\Delta$. Harmonic manifolds are necessarily Einstein manifolds with Ricci curvature $Ric(g)=\kappa g$. Recall that we have the following description of simply connected harmonic manifolds such that $\kappa\geq 0$:
  • (a)
    If $\kappa>0$, then $M$ is a compact rank one symmetric space
  • (b)
    If $\kappa=0$, then $M$ is Euclidean.
Thus simply connected noncompact nonflat harmonic manifolds necessarily have constant negative Ricci curvature $\kappa<0$. For every $v\in SM$, the Busemann function $b_{v}$ satisfies $\Delta b_{v}=h$ (see for instance [Bra24, §2, p. 33]) and so it follows that every harmonic manifold is also asymptotically harmonic. Consequently we obtain the following corollary to Theorem 4.3.
Corollary 4.7.
Let $M$ be a simply connected noncompact (nonflat) harmonic manifold $M$ without focal points. Then the spectrum of the Laplacian is purely absolutely continuous and is contained in $h^{2}/4,\infty)$.
Remark 4.8.
We reiterate here that for this class of manifolds, the negativity of Ricci curvature is not an additional assumption.
We now proceed to give a novel proof that the spectrum is actually equal to the interval $h^{2}/4,\infty)$. For this, we make use of the recently proved [BKP21] unitarity of the Fourier transform for radial functions on harmonic manifolds of purely exponential volume growth, certain estimates on the $c$-function, and the characterisation of harmonic manifolds that the radialisation operator commutes with the Laplacian to pass from radial $\mathrm{L}^{2}$-functions to all $\mathrm{L}^{2}$-functions. We note that unitarity there is shown only in dimensions greater than $5$. However, the only simply connected noncompact (nonflat) harmonic manifolds of dimension less than equal to $5$ are Riemannian symmetric spaces of noncompact type [BKP21, p.145] and this case is covered by the discussion on symmetric spaces at the end.
First, we observe from [Kni16, Theorems 32, 45] that, noncompact harmonic manifolds without focal points fall into the class of noncompact harmonic manifolds of purely exponential volume growth. Let $o$ be an arbitrary but fixed point in $M$. Then for $f\in C_{c}^{\infty}(M)$ which is radial about $o$, the spherical Fourier transform ${\mathcal{F}}^{o}f$ is given by
${\mathcal{F}}^{o}(f)(\lambda,v)=\int\limits_{M}f(y)(\phi_{\lambda}\circ d_{o})(y){\rm{d}}y$
for $\lambda\in{\mathbb{C}}$ and $v\in S_{o}M$. Here $\phi_{\lambda}$ is the unique function on $0,\infty)$ with $\phi_{\lambda}(0)=1$ and satisfying $\Delta(\phi_{\lambda}\circ d_{o})=(\lambda^{2}+\frac{h^{2}}{4})(\phi_{\lambda}\circ d_{o})$, where $d_{o}$ denotes the distance from $o$. Let ${\rm{L}}_{o}^{2}(M)$ denote the closed subspace of ${\rm{L}}^{2}(M)$ consisting of those functions which are radial about $o$ in ${\rm{L}}^{2}(M)$. It is known from [BKP21, Theorem 4.7] that ${\mathcal{F}}^{o}$ is a unitary operator from ${\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$. Here the $c$-function coincides with the Harish Chandra $c$-function on $Im(\lambda)\leq 0$, for rank one symmetric spaces of noncompact type.
We recall that for a real-valued measurable function $f$ on a $\sigma$-finite measure space $(X,\mu)$, the multiplication operator $M_{f}$ on ${\rm{L}}^{2}(X,\mu)$ is self-adjoint with domain $\mathcal{D}(M_{f}):=\{g\in{\rm{L}}^{2}(X,\mu):M_{f}g\in{\rm{L}}^{2}(X,\mu)\}$. Let $m$ be the function on $0,\infty)$ defined by $m(\lambda)=\lambda^{2}+\frac{h^{2}}{4}$. Then we see that the multiplication operator $M_{m}$ on ${\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)$ is self-adjoint on
$\mathcal{D}(M_{m})=\{f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda):M_{m}f\in{\rm{L}}^{2}([0,\infty),C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda)\}.$
Moreover, the spectrum of a multiplication operator $M_{m}$ is equal to the essential range of $m$ [Bor20, Theorem 4.5]. Here
$ess-range(m):=\{x\in{\mathbb{R}}:|m^{-1}(x-\epsilon,x+\epsilon))|>0,\penalty 10000\ \forall\penalty 10000\ \epsilon>0\}.$
where $|\cdot|$ denotes the measure. Now to determine the spectrum of $M_{m}$ we argue as follows. For any $x\in{\mathbb{R}}$ and $\epsilon>0$,
$m^{-1}(x-\epsilon,x+\epsilon)=\{\lambda\in[0,\infty):x-\epsilon<\lambda^{2}+\frac{h^{2}}{4}<x+\epsilon\}$
Note that when $x<\frac{h^{2}}{4}$ and $\epsilon=-(x-\frac{h^{2}}{4})$, $m^{-1}(x-\epsilon,x+\epsilon)$ is empty and so $x\notin ess-range(m)$. Equivalently, $ess-range(m)\subseteq[\frac{h^{2}}{4},\infty)$. Now for any $x\in[\frac{h^{2}}{4},\infty)$ and $\epsilon>0$ we have,
[…]
We recall the estimates for $|c(\lambda)|^{-1}$ given in [BKP21, §4.1].
If $x-\frac{h^{2}}{4}-\epsilon<0$, then $\lambda\in\left[0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)$. If $\sqrt{x-\frac{h^{2}}{4}+\epsilon}\leq K$ then we have
$\frac{1}{C}\lambda\leq|c(\lambda)|^{-1}$
and so
$\int\limits_{m^{-1}(x-\epsilon,x+\epsilon)}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\frac{C_{0}}{C}\int\limits_{0}^{\sqrt{x-\frac{h^{2}}{4}+\epsilon}}\lambda^{2}{\rm{d}}\lambda>0,\quad\text{since}\penalty 10000\ x-\frac{h^{2}}{4}+\epsilon>0$
Now, if $K\in\left(0,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)$ then
$\begin{cases}\frac{1}{C}\lambda\leq|c(\lambda)|^{-1},&\lambda\in[0,K)\\ \frac{1}{C}\lambda^{\frac{n-1}{2}}\leq|c(\lambda)|^{-1},&\lambda\in\left[K,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)\end{cases}$
holds thereby
$\int\limits_{m^{-1}(x-\epsilon,x+\epsilon)}C_{0}|c(\lambda)|^{-2}{\rm{d}}\lambda\geq\begin{cases}\frac{C_{0}}{C}\int\limits_{0}^{K}\lambda^{2}{\rm{d}}\lambda,&\lambda\in[0,K)\\ \\ \frac{C_{0}}{C}\int\limits_{K}^{\sqrt{x-\frac{h^{2}}{4}+\epsilon}}\lambda^{n-1}{\rm{d}}\lambda,&\lambda\in\left[K,\sqrt{x-\frac{h^{2}}{4}+\epsilon}\right)\end{cases}$
and hence the integral is always strictly positive. A similar argument shows that the integral remains strictly positive for the other case in (13). It follows that $ess-range(m)=[\frac{h^{2}}{4},\infty)$ and so the spectrum of $M_{m}$ is equal to $\frac{h^{2}}{4},\infty)$. For functions that are not necessarily radial, a notion of Fourier transform, relative to a point, was introduced in [BKP21, §5] where it was shown to be an isometry but the unitarity is not known. This Fourier transform coincides with the spherical Fourier transform ${\mathcal{F}}^{o}$ for radial functions [BKP21, Proposition 5.3]. Then we see as in [Bra24, Lemma 6.5], that
${\mathcal{F}}^{o}\Delta f=M_{m}{\mathcal{F}}^{o}f,\quad f\in{\rm{L}}_{o}^{2}(M).$
Since ${\mathcal{F}}^{o}$ is unitary and $\mathcal{D}(M_{m})={\mathcal{F}}^{o}(\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M))$, we have the following lemma.
Lemma 4.9.
The spectrum of the Laplacian $\Delta$ on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ is equal to $h^{2}/4,\infty)$.
We now claim that,
Lemma 4.10.
$h^{2}/4,\infty)\subseteq\sigma(\Delta)$ where $\sigma(\Delta)$ is the spectrum of $\Delta$ with domain $\mathcal{D}(\Delta)$.
Proof.
If $\mu$ is not in $\sigma(\Delta)$, then $\Delta-\mu:\mathcal{D}(\Delta)\to{\rm{L}}^{2}(M)$ is invertible with bounded inverse. In particular, $\Delta-\mu$ is injective on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. If $f\in{\rm{L}}_{o}^{2}(M)$, then there exists a unique $g$ in $\mathcal{D}(\Delta)$ such that $(\Delta-\mu)g=f$. We have
$(\Delta-\mu)g=f=R_{o}f=R_{o}(\Delta-\mu)g=(\Delta-\mu)R_{o}g$
and since $\Delta-\mu$ is invertible we have $R_{o}g=g$ i.e., $\Delta-\mu$ is surjective from $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$ onto ${\rm{L}}_{o}^{2}(M)$. Since $(\Delta-\mu)^{-1}$ is bounded on ${\rm{L}}^{2}(M)$ and ${\rm{L}}_{o}^{2}(M)$ is a closed subspace, $(\Delta-\mu)^{-1}$ is also bounded on $\mathcal{D}(\Delta)\cap{\rm{L}}_{o}^{2}(M)$. The claim now follows from Lemma 4.9. ∎
The equality of the spectrum is obtained from Corollary 4.7 and Lemma 4.10. In summary, we have the following proposition.
Proposition 4.11.
Let $M$ be a simply connected noncompact (nonflat) harmonic manifold without focal points. Then the spectrum of Laplacian is purely absolutely continuous and is equal to $h^{2}/4,\infty)$.
As remarked before, rank-one symmetric spaces of noncompact type, Damek-Ricci spaces are among the classes of manifolds which satisfy the hypothesis of Proposition 4.11. In fact, it is known that an irreducible symmetric space is harmonic if and only if it is rank one [Kni16]. We shall see that, on the other hand, Riemannian symmetric spaces of noncompact type and higher rank are not even asymptotically harmonic and therefore Theorem 4.3 does not apply. Nonetheless, we will indicate below that our method can still be employed to show an analogue of Proposition 4.11 in this setting. Let $M$ be an $n$-dimensional symmetric space of noncompact type i.e., $M$ is a simply connected Riemannian manifold with parallel curvature tensor $R$ [Ebe96, Chapter 2]. Note that $M$ is necessarily of the form $G/K$ where $G=NAK$ is the Iwasawa decomposition of a noncompact connected semisimple Lie group $G$ with finite center. The rank of the symmetric space is the dimension $d$ of $\mathfrak{a}$, the Lie algebra of $A$. The Killing form on the Lie algebra of $G$ induces an inner product on $\mathfrak{a}$ denoted by $\langle\cdot$$\cdot\rangle$}. Let us fix a base point $o=eK\in G/K$. If the rank of $M$ is atleast two, then its sectional curvature vanishes on the $G$-conjugates of $A.o$ in $G/K$. These are the maximal totally geodesic flat submanifolds in $G/K$. For $v\in S_{o}M$, consider the geodesic $\gamma_{v}$ satisfying $\gamma_{v}^{{}^{\prime}}(0)=v$ and let $R_{v}$ be the curvature tensor restricted to $\gamma_{v}$. Choose an orthonormal basis $\{e_{1},\cdots,e_{n-1}\}$ of eigenvectors of $R_{v}$ for the subspace in $T_{o}M$ orthogonal to $v$ i.e.,
$R_{v}e_{i}=-R(v,e_{i})v=\lambda_{i}(v)e_{i}.$
From the structure theory for noncompact connected real semisimple Lie algebras, we have the set of all positive roots denoted by $\sum^{+}$. Let $m_{\alpha}$ denote the multiplicity of $\alpha\in\sum^{+}$. The finite reflection group $W$ generated by the roots is known as the Weyl group. We denote the tangent vector that is dual to $2\rho=\sum\limits_{\alpha\in\sum^{+}}m_{\alpha}\alpha$ also by $2\rho$. The arguments in [Wan15, Lemma 1] show that for each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function given by $\sum\limits_{i}\sqrt{\lambda_{i}(v)}={\langle v},{2\rho\rangle}$. This constant is the mean curvature of horospheres with respect to $b_{v}$. Here we argue as in Proposition 4.4 by using Stoke’s theorem to get a lower bound for Cheeger’s constant namely, ${\langle v},{2\rho\rangle}\leq h_{Cheeger}$ for all $v\in S_{o}M$. In particular, for $v=\frac{\rho}{\left\lVert\rho\right\rVert}$ we have $\left\lVert\rho\right\rVert\leq h_{Cheeger}$ and hence by Cheeger’s inequality we obtain
(14) $\left\lVert\rho\right\rVert^{2}\leq\lambda_{0}(M).$
Observe from [Wan15], that along a geodesic $\gamma_{v}$ the Jacobi field $Y_{i}$ satisfying the initial condition $Y_{i}(0)=0$ and $Y_{i}^{{}^{\prime}}(0)=e_{i}$ is given by the explicit formula
$Y_{i}(t)=\frac{1}{\sqrt{\lambda_{i}(v)}}\sinh\left(\sqrt{\lambda_{i}(v)}t\right)E_{i}(t)$
where $E_{i}$ is a parallel orthonormal frame field along $\gamma_{v}$ with $E_{i}(0)=e_{i}$. Then we see that
$Y(t)=\sum\limits_{i=1}^{n}{\langle w},{e_{i}\rangle}Y_{i}(t)$
solves the Jacobi equation (1) with initial conditions $Y(0)=0$ and $Y^{{}^{\prime}}(0)=w$ for arbitrary $w\in SM$. That $\left\lVert Y(t)\right\rVert$ is strictly increasing can then be deduced from the fact that $\sinh$ is an increasing function and so, it follows from [O’S74, Proposition 4] that $M$ has no focal points. It is known that $M$ is also Einstein with negative Ricci curvature [Ebe96, §2.14, p. 96]. Therefore Theorem 4.1 applies and we have,
  1. (i)
    $\left\lVert\nabla F\right\rVert\leq 1$ since $\left\lVert\nabla b_{v}\right\rVert=1$ for all $v\in S_{o}M$.
  2. (ii)
    From $\Delta b_{v}={\langle v},{2\rho\rangle}$ for all $v\in S_{o}M$, we have $|\Delta b_{v}|\leq\left\lVert 2\rho\right\rVert$. Thus $\Delta b_{v}$ is bounded independent of $v\in S_{o}M$ by $\left\lVert 2\rho\right\rVert$ and so $|\Delta F|\leq\left\lVert 2\rho\right\rVert$.
  3. (iii)
    For each $v\in S_{o}M$, $\Delta b_{v}$ is the constant function ${\langle v},{2\rho\rangle}$ and so we have $\Delta^{2}b_{v}=0$. It follows that, $\Delta^{2}F=0$.
As before, the absolute continuity of the spectrum $\sigma(\Delta)$ is established by using [Xav88, Theorem 2]. That the spectrum is equal to $\left\lVert\rho^{2}\right\rVert,\infty)$ can be shown analogous to the case of harmonic manifolds by using the spherical Fourier transform and estimates on the c-function found in [GV88, Chapter 3, p. 101, Theorem 6.3.4].
5. Funding
This work was supported by the Centre for Operator Algebras, Geometry, Matter and Spacetime, Ministry of Education, Government of India through Indian Institute of Technology Madras [SB22231267MAETWO008573 to A.P. and S.B.]; and by the National Board for Higher Mathematics, Department of Atomic Energy, Government of India [Ref.No:0203/11/2019-R&D-II to S.B.]